A Bibliography of the Annual IEEE Hot Chips Symposia

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2024 Version 0.36

Title word cross-reference

16 [ABG^{+}16]. 16 × 16 [AEJ^{+}00]. 28 [KBN16]. 3 [Awa95, Bat97, DFG^{+}13, Jay98, Kre98, LD98, MMG^{+}98, Pot97, Tre96, Tre97]. 64 × 64 [Wei00].

-nm [ABG^{+}16, KBN16].

0.35-micron [BB96]. 0.5W [San96]. 000-Processor [BSP^{+}17].

1 [BH15, Bre10, Kru00, Ste95]. 1/4 [PAGC^{+}97]. 1/4-Inch [PAGC^{+}97]. 100 [PSW91, Pot97]. 100-Mops [PSW91]. 1000-Way [LL98]. 100Kbit [Oru94]. 100Kbit/s [Oru94]. 110 [San96]. 1100 [Sla97]. 115W [Ano93a]. 12 [DTB01].

12-bit [OKN^{+}00]. 13 [KW02]. 1300 [SLR^{+}99]. 14 [WD03]. 15 [FD04]. 150 [SHMS95]. 1500 [Gan98]. 16 [DD05, FH99]. 16-Core [FJL^{+}13]. 16-nm [FME18]. 160MHz [San96]. 16bps [CEF^{+}99]. 17 [SS06]. 18 [KS07]. 19 [AM08]. 19-20 [Rei96]. 196 [CES^{+}11]. 1993 [IEE93]. 1994 [IEE94]. 1995 [IEE95]. 1996 [IEE96, Rei96]. 1997 [IEE97]. 1999 [IEE99]. 1GHz [MBB^{+}99].

2 [KSI^{+}96, Lee97, MS03, NTK^{+}97, Nga95, OWJF98, RMC04, SSB20]. 2-Petaflop [SB23]. 2.0 [Lee97, VBC^{+}21]. 2000 [IEE00]. 2001 [IEE01]. 2002 [IEE02]. 2003 [IEE03]. 2004 [IEE04]. 2005 [IEE05]. 2006 [IEE06]. 2007 [IEE07]. 2008 [IEE08]. 2009 [IEE09]. 2010 [IEE10]. 2011 [IEE11]. 2012 [IEE12]. 2013 [IEE13]. 2023 [LV24]. 21 [AW10]. 21164 [BB96, ERPR95]. 21264 [Kes98].

21264a [BVD^{+}99]. 215Hz [CVS^{+}00]. 21st [Sam99]. 22 [RE11]. 23 [BB12]. 230MHz [O’D99]. 24 [KZ13]. 25 [NN14]. 26 [NS15]. 27 [AR16a]. 28 [CM17]. 28-nm [CCA^{+}19]. 29 [Eec18]. 29K [McM95].

3 [HWG^{+}09, Tre98]. 30 [KR19, SB23]. 30-Teraflops [SB23]. 30-Teraflops/W [SB23]. 300 [Ano93a, Gan98, KS90, Nt97]. 3171 [BSC^{+}90]. 32 [CHH^{+}98, KS90, Kur21, RY21, Rub97, RDJ^{+}13]. 32-bit [NTK^{+}97, KS90]. 32-nm [RDJ^{+}13]. 32-Way [KAO05]. 32b [San96, Ano93a]. 33 [Joh22, SS22]. 34 [DA23, Joh23]. 3DNow [OWJF98].

4 [GDES08]. 4-Gbps [GDES08]. 4-Inch [PAGC^{+}97]. 40-nm [Man09]. 400 [DRM^{+}23]. 400-G [DRM^{+}23]. 4096-Core [ZSB21]. 440LX [Mal97]. 4MB [Shi98]. 4th [BT24]. 4x [KKK^{+}99].

5 [Bro00]. 5-qubit [CVS^{+}00]. 5.5 [KIS^{+}99]. 50Mpixel [OKN^{+}00]. 50Mpixel/s [OKN^{+}00]. 511-Core [DXT^{+}18]. 5W [Bur97, O’D99].

600 [Kes98, LL98]. 6000 [Ari96, OB91]. 64 [Fan99a, Kni99a, Kni99b]. 64-bit [BBTV15, LL98, Naa95, Nt97, She95, SBKK99, Tre95]. 64-Core [DFG^{+}13]. 65nm [DAV06]. 6800 [MM05]. 6M [RMC04]. 6th [DKyL^{+}17]. 6th-Generation [DKyL^{+}17].

7040 [SKP24].

8000 [Naa95]. 8500 [Joh98]. 870 [BCC^{+}02].

90nm [FSP06].

A/V [GDES08]. A100 [CGG^{+}21]. Academia [Eec17]. Accelerated [BCF^{+}14, KBN16]. Accelerating [Bir98, DDC^{+}98, ESG^{+}05, KKSS99, Lee95, TKS^{+}99]. Acceleration [GHY^{+}17, SB23]. Accelerator [Buc97, DXT^{+}18, Kre98, MMG^{+}98, PAA^{+}06, Pia98, YYA06, Pri90, Dja96]. Access [O’C00a]. Across [DDC^{+}98]. Active [PAGC^{+}97]. Adapter [Edd02]. Adaptive [DRM^{+}23, FAWR^{+}11]. Address [Kut99, O’C00b, O’C00a, OG01, Vit00]. Advanced [SF18]. Advances [Hun97]. Afternoon [Dub97, Pra96]. Age [DPY18]. Agile [LWC^{+}16]. AGP [KKK^{+}99]. AGPset [Mal97]. Ahead [Var94]. AI [MD20]. Alchemy [Plu00a]. Algorithm [CD95]. Algorithms [Vis99]. All-Programmable [FME18]. Alpha [Ano93b, BVD^{+}99, BB96, ERPR95, Kes98, Rub97]. AltiVec [DDC^{+}98, Phi98]. Alto [IEE98]. Always [BBC^{+}15]. Always-on [BBC^{+}15]. AMD [BT24, BCF^{+}14, BFS12, BCD^{+}11, Chr95, Chr96, CKD^{+}10, DRM^{+}23, KKK^{+}99, KMAC03, OWJF98, OS08, SKP24, SSB20]. AMD-K5 [Chr95, Chr96]. AMULET2e [Gar96]. AMULET3i [Gar00]. Analog [OKN^{+}00, OW01]. Anatomy [THT^{+}04]. Annual [Eec16, Ste90a, Ste90b]. Appliances [JSR^{+}99]. Application [GHSV^{+}11]. Applications [BYM^{+}06, BBC^{+}15, BSP^{+}17, FM95, HYYS96, KTP^{+}99, LCP^{+}11, Rub97, SC91, SF18, TSI06, Ano93b, Yea96]. Applied [RSS98]. Approach [BBSG11, KR96, LWC^{+}16]. APU [BFS12]. Architecting [CM00]. Architectural [Bro00, Dub97]. Architecture [Als90, Ano93b, Cas00, CEM^{+}95, CAV^{+}14, CH06, DPY18, GM21, Gol00, GHSV^{+}11, GA21, Hed00, Joy96, Kni99a, KFL99, KTP^{+}99, KIS^{+}00, LD98, Lie23, LNOM08, Nem95, Nga95, OS08, Phi98, PSS^{+}91, Rat98, RNA^{+}12, STKS17, SL00, TUHwH99, Tre99, Tru97, WKP11, Yeh06, YYA06, ZSB21, Chr96, Hes07, OB91, Pri90, SMHB91, TO96a, BDH03]. Architectures [DXT^{+}18, MD20, Vis99]. Arena [Joh20]. ARM [BBTV15, San96, SBB^{+}17, GBW^{+}23, PSB^{+}20, SSR21]. Arm-Based [SSR21]. ARM810 [Lar96]. ARM9E [Seg99]. ARM9ESP [Bur99]. ARM9TDMI [Bur99]. Array [ABG^{+}16, BYM^{+}06, BSP^{+}17]. AsAP [BYM^{+}06]. ASIC [CC95, Man09, Pfi99]. associated [Mal97]. Asynchronous [BYM^{+}06, Gar00]. At-Memory [SB23]. Athlon [KKK^{+}99]. Atom [BvdGM^{+}15]. ATSC [Par98]. Attached [Gan98]. Au1000 [Plu00a]. Audio [FM95, Sav98, Ste95]. Audio/Video [Ste95]. Auditorium [IEE98, IEE13]. August [IEE93, IEE94, IEE95, IEE96, IEE97, IEE98, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, IEE12, IEE13, Rei96]. Automatically [AAW^{+}96]. Automotive [SF18]. AV [ASK97, SANK98]. Availability [Qua00]. AXP [Ano93b]. Azure [Sti19].

Bandwidth [SL00, WAA^{+}20]. Based [GBW^{+}23, SSR21, WMSH09, IDTS00]. Basics [Kni99a]. Basics/Introduction [Kni99a]. BCM4100 [FH99]. BCM4100/BCM4210 [FH99]. BCM4210 [FH99]. BCM5600 [EM99]. Beast [Gar95]. Beat [Lar96]. Behind [Sti19]. Below [FSP06]. Benchmark [MRC^{+}20, AAW^{+}96]. Berkeley [CFK^{+}10]. Best [Bas00, RY21, WBC^{+}96]. Better [Ber98, Gar95]. Beyond [Hes07, LCP^{+}11]. Big [MMG^{+}98]. Binary [CHH^{+}98]. bipolar [Ano93a]. bit [BBTV15, LL98, Naa95, NTK^{+}97, Nt97, OKN^{+}00, She95, SBKK99, Tre95, KS90]. Bitcoin [BH15]. Blitzen [Kre98]. Block [BCC^{+}02]. Blocking [AEJ^{+}00]. Blocks [PSB^{+}20]. Blue [HOF^{+}12]. board [MKN^{+}98]. Bobcat [BCD^{+}11]. Boost [Gol00]. bottleneck [Joh90]. Boundaries [NCT^{+}98]. Brain [SKW^{+}23]. Brainwave [CFO^{+}18]. bridge [WBC^{+}95, PKB^{+}15, RNA^{+}12]. bridge/memory [WBC^{+}95]. Brief [Bar21]. Bringing [Khu96]. Broadband [Gol00, Gre11, Sam99, Sam00b]. Broadcasting [Hun97]. Broadcom [EM99, SP09]. BROOM [CCA^{+}19]. Building [BCC^{+}02, Gar95, LWC^{+}16, PSB^{+}20, SP09]. Built [MD20]. Bulldozer [BBSG11]. Buses [Jam90].

C [Bro00]. C-5 [Bro00]. C-Port [Bro00]. C2000 [BvdGM^{+}15]. C400 [SMHB91]. C64x [Gol00]. CA [IEE93, IEE94]. CA1024 [SBS^{+}06]. Cache [Bur97, CKD^{+}10, Faa98, RMC04]. Caches [Cha96, VJFG17]. California [IEE95, IEE96, IEE97, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, IEE12, IEE13, IEE98]. Calisto [NIJ^{+}03]. Camera [Fos98]. Can [Ano03]. Capabilities [vES98]. Capacity [Shi98]. Carrizo [KBN16]. Cartridge [Sam00a]. Cascade [AFK^{+}19]. Case [PAY96]. Casting [Pfi99]. CD [FM95]. CDMA [She99b]. Celerity [DXT^{+}18]. Center [IEE12]. Centip3De [DFG^{+}13]. Century [Sam99]. Cerebras [Lie23]. Chairs [JW98]. Challenge [Wha98]. Challenges [Pen90, Rab06, Won03, Mal97]. Change [Gon99]. Channel [Edd02]. CHERI [GBW^{+}23]. CHERI-Based [GBW^{+}23]. Chess [hH98, hH98]. Chip [ABG^{+}16, AEJ^{+}00, Ari96, Awa95, ASN^{+}99, BCF^{+}14, BWBJ11, Bur97, CD95, DRM^{+}23, EGL^{+}90a, EM99, FM95, FAWR^{+}11, Fos98, FH00, Gar00, HOF^{+}12, hH98, Joh20, KST04, KML04, Kec97, KSIA95, McC99, NIJ^{+}03, NCT^{+}98, Oru94, PAGC^{+}97, Pet00, Plu00a, Pot97, Rat98, SC91, SO14, SGG^{+}12, Shi98, Ste95, SBS^{+}06, TSW^{+}01, Wei00, Ano93d, KSI^{+}96, MKN^{+}98, TO96a]. Chiplet [WAA^{+}20, ZSB21]. Chips [AS95, Alt13, Alt14, AAFH95, AM08, AR16a, AR16b, Ano95, Ano00, Ano03, AW10, BS98, BB12, CM17, DTB01, DD05, DXT^{+}18, DA23, Dit00, Eec15, Eec16, Eec17, Eec18, FD04, HW91, IEE94, IEE97, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, IEE12, IEE13, Joh19, Joh20, Joh22, Joh23, Joh90, KZ13, KB20, KW02, KS07, KR19, Kur21, LV24, Mat97, NN14, NS15, NPY^{+}21, RY21, Rat06, Rei96, RE11, RC13, SS22, SS06, Ste90a, Ste90b, WD03, IEE93, IEE95, IEE96, IEE98, JA96, Alt11, Hoo90, Jou92, KvdW09, Var94]. Chips-III [Jou92]. Chipset [CEF^{+}99, FH99, GDES08]. Chipsets [Par98]. Choices [Ano95]. Circuit [Kid14]. Circuits [TKM^{+}02]. Classifier [IDTS00]. ClassiPI [IDTS00]. Clipper [SMHB91]. clock [Cra90]. Clockless [Cum04]. close [hH98]. Cloud [PSB^{+}20]. Cloud-to-Edge [PSB^{+}20]. Clouds [MFN^{+}17]. Cluster [BDH03]. CMOS [San96, AEJ^{+}00, Ano94c, CCA^{+}19, Faa98, PAGC^{+}97, RDJ^{+}13]. CMP [CH06, HHS^{+}99, HHS^{+}00]. CMT [CCE^{+}09]. Co [Hay97, JW98, KKO06, Lie23, SKW^{+}23]. Co-Chairs [JW98]. Co-Design [Lie23]. Co-Designed [SKW^{+}23]. Co-Processor [Hay97, KKO06]. Code [DKyL^{+}17, RNA^{+}12]. Code-Named [DKyL^{+}17, RNA^{+}12]. Codes [Rat06]. Codesign [GHY^{+}17]. CoinTerra [BH15]. Collaborative [Mey06]. Color [BD99]. Combining [TO96a]. Commercial [SBKK99]. Commodity [Ros99]. Common [Man09]. Communications [CAV^{+}14, FME18, Gol00, Hun97, LS98, NIJ^{+}03, Sam99, She99a, She99b, Sla97]. Companies [Bas00]. Compatible [Bos96]. Compilation [Fan99b]. Compiler [ADG^{+}96, Fan99a, Pen90, TGK^{+}96]. Compilers [AAFH95, KFL99]. Complexity [MM96]. Compliant [Par98]. Compositing [Dja96]. Compression [AHM^{+}00, CD95, Nga95]. Computation [SVC01, CVS^{+}00]. Computational [TKM^{+}02]. Compute [BBSG11, HOF^{+}12, TSV^{+}20, VBC^{+}21]. Computer [Bre10, DPY18, Kut99, RSS98, SKW^{+}23, TSV^{+}20, TSW^{+}23]. Computing [Bar21, BJ06, CFK^{+}10, CSM^{+}21, DHM97, Dit00, KKSS99, LNOM08, LCP^{+}11, MYK^{+}10, McC99, ND10, TKS^{+}99, YHT^{+}15, ZSB21]. concurrency [Yea96]. conference [IEE98]. Configurable [DHM97, Gon99, Gon00]. Configuration [MKN^{+}98]. Confronting [Wha98]. Connected [Sam99]. Connecting [FH00]. connections [SL00]. Considerations [Wei96]. Consoles [ML21]. Consumer [FM95, KTP^{+}99]. Content [IDTS00]. Continue [Jam90]. Continuous [ABD^{+}97]. Controller [Bur97, NABR95, TSW^{+}01, Tre98, WBC^{+}95]. Converged [PKB^{+}15]. Convey [Bre10]. Cool [Ano03, Dit00, Rat06, Ano93d]. Cooler [Bal95]. Coprocessor [DKB^{+}90, Bar97]. Coprocessors [BSC^{+}90, CSM^{+}21]. Core [BT24, CGG^{+}21, CC95, DXT^{+}18, DKyL^{+}17, DFG^{+}13, FZW^{+}12, FJL^{+}13, HMB^{+}14, Hes07, KST04, Kru00, MB05, Sha00a, Sha00b, ZSB21]. Cores [CSM^{+}21, SB23]. Cost [BCC^{+}02, Luc99, SBS^{+}06, Ano93b, KSI^{+}96, SLR^{+}99]. Cost-Effective [BCC^{+}02, SBS^{+}06, KSI^{+}96]. Coupled [LD98]. Court [WBC^{+}96]. CPU [Cra90, Gan98, HMR96, Kum96, Mod97, Nt97]. CPUs [Ber98, BT24, ESG^{+}05, Kur21]. Crossbar [Cum04, Wei00]. Crusoe [Dit00, Fle00]. Cryptocurrency [BH15]. Cryptography [Bir98]. Cryptosystems [ESG^{+}05]. Crystal [BD99]. Cubes [Ano03]. Cupertino [IEE12]. Custom [Dja96, Faa98]. Cutting [Eec17, Fle00, LB00]. Cutting-Edge [Eec17, LB00]. CW4010 [CC95]. Cycle [Pra96, Cra90]. Cycles [ABD^{+}97].

D [Awa95, Bat97, DFG^{+}13, Jay98, Kre98, LD98, MMG^{+}98, Pot97, Shi98, Tre96, Tre97]. DAC [Dja96]. Dancing [Lar96]. Dark [GHSV^{+}11]. Datacenter [BvdGM^{+}15, CFO^{+}18]. DataPlay [Dav02]. Datawave [SC91]. Day [Ano03]. DDC [Kid14]. Debuts [AHM^{+}00]. DECchip21066 [Ano93b]. Decoder [Ste95]. Decoding [MD06]. Deep [Lie23]. Defining [War97]. Definition [MD06]. Delay [NTK^{+}97]. Delta [Tre96]. dense [FSP06]. Denver [BBTV15]. Design [BTR02, BB96, DXT^{+}18, Dub97, EGL^{+}90b, EGL^{+}90a, Gon99, Joh20, Lie23, MBB^{+}99, NPY^{+}21, PKB^{+}15, RSS^{+}08, SMHB91, SBKK99, WP97, Won03, Joh90, Pap96]. Designed [SKW^{+}23]. Designing [CSM^{+}21, hH98, WBC^{+}95]. Designs [LB00]. Desktop [Khu96]. Developing [BSC^{+}90, Chr96, Pri90]. Development [Mey06, Chr96, Mal97]. Device [DHM97, SB23]. Devices [Vit00]. Did [Joh20, hH98]. dies [Pap96]. Different [Lar96]. Digital [FME18, Fos98, OKN^{+}00, OW01, PAGC^{+}97, Sav98, TP10, THT^{+}04, Rub97]. Digital-RF [FME18]. Directed [CHH^{+}98]. Direction [Gre11]. DirectX [Tre98]. Discussion [vdWAB^{+}06, GTB99]. Display [BD99]. Distributed [GM21, NABR95]. Distribution [Dav02, DHM97]. Dive [Lie23]. Diverse [Eec15]. DNNs [CFO^{+}18]. Do [ABD^{+}97]. DOJO [TSW^{+}23]. DRAM [KGM^{+}00, LD98, O’C00b, O’C00a, PAY96, Shi98]. DRAMs [Prz97]. Driven [DSK^{+}92]. Driving [TSV^{+}20]. Drum [Lar96]. DSP [CAV^{+}14]. DTV [Par98, Rat98]. Dual [KST04, MB05]. Dual-Core [KST04, MB05]. Dual-Thread [MB05]. Dust [WAP00]. Dynamic [Fan99b, Mod97]. Dynamically [SGG^{+}12, YYA06].

ECL [Ano93a, BAC^{+}90]. Economics [WD03]. Edge [BWBJ11, CSM^{+}21, Eec17, LB00, PSB^{+}20, Plu00a]. Edge-Computing [CSM^{+}21]. Editors [AS95, AM08, AW10, BS98, DTB01, FD04, HW91, JA96, KW02, KS07, LB00, SS06, WD03]. Effective [BCC^{+}02, SBS^{+}06, KSI^{+}96]. Effectiveness [Lee97]. Effects [Joh98]. Efficient [Bat97, BvdGM^{+}15, DSK^{+}92, FZW^{+}12, GHY^{+}17, KBN16, MD06, TUHwH99]. efficiently [Yea96]. Eight [FJL^{+}13]. Electronics [RSS98]. eliminate [Joh90]. Embedded [ASK97, Cum04, KGM^{+}00, LD98, O’C00b, O’C00a, SANK98]. Emerging [Joh19]. Emotion [KIS^{+}99, KIS^{+}00]. Emphasizing [Yea96]. Empowering [DPY18]. EMU10K1 [Sav98]. Emulation [HWG^{+}09]. Emulator [HMR96]. Enabling [Sam99, Seg99, Vit00]. Encoder [KSIA95, MKN^{+}98, Nga95, KSI^{+}96]. End [OKN^{+}00, OW01, Vin07]. Endian [Jam90]. Energy [FAWR^{+}11, KBN16]. Energy-Efficient [KBN16]. Engine [ACD^{+}00, Sel18]. Engines [NABR95]. Enhanced [Luc99, SLR^{+}99, KGM^{+}00, Lee95]. Entertainment [Kut99, KKO06]. entertainment-quality [KKO06]. EPI41100 [CEF^{+}99]. EPI41210 [CEF^{+}99]. EPI41210/EPI41100 [CEF^{+}99]. Epigram [CEF^{+}99]. EPYC [BT24]. Era [ND10]. Establish [NMP^{+}96]. Established [Bas00]. Estimation [KSIA95]. Ethernet [AEJ^{+}00, EM99, MD20]. Evaluation [EG95, GBW^{+}23]. Evening [WBC^{+}96]. Evolving [Bal95, Hes07]. Exa [TSW^{+}23]. Exa-Scale [TSW^{+}23]. executing [Cra90]. Execution [EG95, Kes98, Mod97, Rub97, ERPR95]. Expanding [NCT^{+}98]. experience [KKO06]. Exploiting [Alt13]. Exploring [FZW^{+}12]. Exponentiation [Oru94]. Express [CRTI00]. Extensible [Gon99, Gon00]. Extension [SBB^{+}17, TUHwH99]. Extensions [Gol00, Lee97, Mah96, Tha99].

Fabric [BJ06, DXT^{+}18, TKM^{+}02]. Fabrics [Wei00]. Face [WD03]. Facing [KML04]. Families [Bur99]. Family [Als90, Bal95, BvdGM^{+}15, McM95, OS08, Plu00a, Seg99, SL00, Yeh06]. Fast [Ber98, CD95, DXT^{+}18, MMG^{+}98, O’C00b, O’C00a, OW01, Rub97]. Faster [Bal95]. Fat [VJFG17]. Fault [RSS^{+}08]. Fault-Tolerant [RSS^{+}08]. Feast [Eec16]. Feature [SHMS95]. Features [FAWR^{+}11, Kni99b, Naa95, Qua00]. Fermi [WKP11]. Fi [FM95]. Field [ABG^{+}16, BD99]. Field-Programmable [ABG^{+}16]. Field-Sequential [BD99]. final [Pap96]. Fine [BSP^{+}17]. Fine-Grained [BSP^{+}17]. First [BH15, BBTV15, Kag96, Lie23, McM95, Plu00a, Ste90a, Ste90b]. First-Generation [BH15]. Five [SVC01]. Five-Qubit [SVC01]. Flexibility [SL00]. Flint [IEE12]. Floating [BSC^{+}90, DKB^{+}90, ZSB21]. Floating-Point [BSC^{+}90, DKB^{+}90, ZSB21]. Flying [Chr96]. Forum [ES99, GTB99]. Forward [Joy96]. Forwarding [ACD^{+}00, O’C00b, O’C00a]. four [TO96a]. four-issue [TO96a]. Fourth [HMB^{+}14]. Fourth-Generation [HMB^{+}14]. FPGA [DAV06, Man09]. Frame [Nga95]. Frequency [RMC04, SBJ13]. Fresh [KR96]. Front [OKN^{+}00, OW01]. Front-End [OW01]. Fujitsu [YHT^{+}15]. Full [PAGC^{+}97, TSV^{+}20]. Fully [SBS^{+}06]. Functions [PAGC^{+}97]. Fusion [BFS12]. Future [AAFH95, CH06, GHSV^{+}11]. FUZION [McC99]. FX [Rub97, CHH^{+}98].

G [DRM^{+}23]. G5 [SAC^{+}98]. Game [ML21]. Gamma [Tre97]. Gate [ABG^{+}16]. Gaudi [MD20]. Gbps [GDES08]. GeForce [MM05]. Gen [BT24, PSB^{+}20]. Gene [HOF^{+}12]. Gene/Q [HOF^{+}12]. General [ESG^{+}05, TKM^{+}02]. General-Purpose [ESG^{+}05, TKM^{+}02]. Generation [AJK^{+}15, AFK^{+}19, BH15, BT24, Bir98, DAV06, DKyL^{+}17, ESG^{+}05, HMB^{+}14, KSSF10, KKK^{+}99, Mah96, MYK^{+}10, Phi98, She95, SBJ13, SGC^{+}16, SSR21, Tre96, Tre98, Vit00, Web08, YHT^{+}15, IDTS00]. Genesis [Cho98]. Geometry [Kre98, TUHwH99, Tre97]. Geoscience [LCP^{+}11]. GF100 [WKP11]. GFLOPS [KIS^{+}99]. Gigabit [AEJ^{+}00]. Gigascale [Mei97]. GLINT [Tre96, Tre97]. Gmicro [KS90]. Gmicro/300 [KS90]. goals [Pap96]. Godson [FZW^{+}12, HWG^{+}09]. Godson-3 [HWG^{+}09]. Godson-T [FZW^{+}12]. Golden [DPY18]. Goldstrike [BH15]. Good [Ber98, Joh90]. Google [BDH03, NPY^{+}21]. GPS [KTP^{+}99]. GPU [Bur20, CGG^{+}21, Cho23, FD17, ND10, WKP11]. GPUs [Kur21]. Grained [BSP^{+}17]. Grandmaster [hH98]. Graphics [Awa95, Eer97, Jay98, Khu96, KBN16, LNOM08, MMG^{+}98, Pia98, Tre96, Tre98, Ano93c, Pri90, MM96]. GreenDroid [GHSV^{+}11]. Griffin [OS08]. Guest [AS95, AM08, AW10, BS98, DTB01, FD04, HW91, JA96, KW02, KS07, LB00, SS06, WD03]. GX [Pri90].

H100 [Cho23]. Habana [MD20]. HAL [She95, EG95]. HALO [SKW^{+}23]. Hardware [BYM^{+}06, BVD^{+}99, Bir98, BJ06, GHY^{+}17, Kal96, Lie23, SKW^{+}23, Sti19, YYA06, TO96b]. Hardware/Software [Kal96, Lie23]. Haswell [HMB^{+}14]. HC [Bre10]. HC-1 [Bre10]. HD [GDES08]. HDD [TSI06]. HDTV [Hun97, SBS^{+}06]. Hearing [WMSH09]. Heart [Ari96]. Heat [Joh20]. heavily [Ano95]. Heterogeneous [MD06]. Hexagon [CAV^{+}14]. Hi [FM95]. Hi-Fi [FM95]. Hidden [ML21]. hiding [Yea96]. Hierarchy [CKD^{+}10]. High [ASN^{+}99, Bat97, BTR02, Cha96, CCE^{+}09, CM00, Cum04, GBW^{+}23, HBG^{+}97, Jay98, Kru00, Kum96, LCP^{+}11, Luc99, Mod97, MD06, O’C00a, Plu00a, Plu00b, Pot97, Qua00, RC13, Shi98, SBJ13, SLR^{+}99, TP10, WAA^{+}20, Yeh06, YHT^{+}15, Ano94a, IEE93, IEE94, TO96a]. High-Bandwidth [WAA^{+}20]. High-Frequency [SBJ13]. High-Performance [ASN^{+}99, CCE^{+}09, CM00, Cum04, GBW^{+}23, HBG^{+}97, Jay98, Kru00, LCP^{+}11, Yeh06, YHT^{+}15, IEE93, TO96a]. High-Speed [O’C00a, TP10, SLR^{+}99]. Higher [RMC04]. highest [AAW^{+}96]. Highlights [AR16b]. Highly [CD95, EM99, Nt97, Pro06, Ano93c]. highly-integrated [Ano93c]. History [Bar21]. HL [MKN^{+}98]. Home [CEF^{+}99, FH00, JSR^{+}99, KKO06, Rab06]. Hopper [Cho23]. Horus [KO05]. Hot [Alt13, Alt14, AAFH95, AR16b, Ano95, Ano00, Ano03, BS98, Eec15, Eec16, Eec17, HW91, Joh19, Joh90, JA96, KB20, Rat06, Ste90a, Ste90b, Var94, Ano93d, AS95, Alt11, AM08, AR16a, AW10, BB12, CM17, DTB01, DD05, DA23, Eec18, FD04, Hoo90, IEE93, IEE94, IEE95, IEE96, IEE97, IEE98, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, IEE12, IEE13, Joh22, Joh23, Jou92, KvdW09, KZ13, KW02, KS07, KR19, Kur21, LV24, Mat97, NN14, NS15, RY21, Rei96, RE11, SS22, SS06, WD03]. HP [Kum96]. Human [WMSH09]. Hybrid [Pro06]. Hydra [HHS^{+}99, HHS^{+}00]. Hyperthreading [KM03].

I-Frame [Nga95]. I/O [WAA^{+}20, Ber98]. i486 [Cra90]. i740 [Pia98]. IA [Fan99a, Fan99b, Kni99a, Kni99b, RDJ^{+}13]. IA-32 [RDJ^{+}13]. IA-64 [Fan99a, Kni99a, Kni99b, Fan99b]. IA64 [KFL99]. IBM [Ari96, AHM^{+}00, BWBJ11, HOF^{+}12, KST04, KSSF10, OB91, RSS^{+}08, STKS17, SBJ13, SAC^{+}98, STSM21, TSW^{+}01, Web08]. IBMLZ1 [CD95]. IC [Bos96, BJ06]. IC’s [Sam99]. If [War97]. iFlow [O’C00b, OG01]. II [Buc97, Mod97, Hes07, HW91, KTP^{+}99, She99b]. III [AHM^{+}00, Jou92, LL98, Nt97, NCT^{+}98]. Illinois [CFK^{+}10]. Image [KDR^{+}00]. IMAGINE [KDR^{+}00, KDK^{+}01]. Imaging [Gol00, OKN^{+}00, OW01]. Impact [Won03]. Implementation [Bat97, EGL^{+}90b, HBG^{+}97, Kag96, Kru00, TO96b, YYA06, SMHB91]. Implementing [BAC^{+}90, Gar95]. Implications [Dub97]. In-Package [WAA^{+}20]. Inc. [Plu00a]. Inch [PAGC^{+}97]. Incomplete [Alt13]. Increasingly [Eec15]. independent [Chr96]. Industries [RSS98]. Industry [Eec17, MRC^{+}20]. Inference [MD20, SB23]. Inferno [WP97]. InfinBand [Cas00]. InfiniBand [Ano00, Edd02]. InfiniBridge [Edd02]. InfiniteReality [MM96]. Information [Dav02, JSR^{+}99, Mey06]. Infrastructure [PSB^{+}20]. initial [Pap96]. Innovate [Bas00]. Innovation [CGG^{+}21, Mey06, Seg99, WD03]. Innovations [Bre10]. Innovative [LD98]. Insensitive [NTK^{+}97]. Instruction [BVD^{+}99, Bre10, DHM97, Mah96, vES98, ERPR95]. Instructions [Pra96, Cra90, TO96a]. Instructions/Cycle [Pra96]. instruments [Chr96]. Integrated [Edd02, EM99, Jay98, Nem95, Nt97, NABR95, Pro06, Ano93c]. Integration [Mei97, Pet00]. Intel [AFK^{+}19, BCC^{+}02, BvdGM^{+}15, DKyL^{+}17, HMB^{+}14, Mal97, Pia98, RNA^{+}12, Sha00a, SGC^{+}16, Wei96]. Intellectual [RSS98]. Intelligent [PAY96]. Interaction [Kal96]. Interconnect [FD17]. Interconnects [Ano00]. Interface [ASK97, FCD^{+}99, PAGC^{+}97]. Interfaces [SF18, SKW^{+}23]. Internal [Shi98]. Internet [Plu00a, Plu00b, Tha99]. Introducing [FAWR^{+}11]. Introduction [AS95, AM08, AW10, BS98, DTB01, ES99, FD04, HW91, Jou92, Kni99a, KW02, KS07, LB00, SS06, WD03, JA96]. IP [ACD^{+}00].

IP/MPLS [ACD^{+}00]. IQ2000 [SL00]. IRAM [KGM^{+}00, PAY96]. Issue [DA23, LV24, SS22, TO96a]. Itanium [MS03, MB05, Qua00, RMC04, Sam00a, Sha00a, Sha00b]. Ivy [PKB^{+}15]. Iwarp [PSW91]. IX [IEE97, Mat97].

Java [Sha96, HBG^{+}97, TO96b, WBC^{+}96]. Job [Alt13].

K5 [Chr95, Chr96]. K6 [OWJF98]. K6-2 [OWJF98]. Kabini [BCF^{+}14]. Key [Bir98, ESG^{+}05]. Keynote [Hes07, Kut99, Pap98, Sam99, Vin07, Vit00]. KiloCore [BSP^{+}17]. Kinect [SO14]. Klessydra [CSM^{+}21]. Klessydra-T [CSM^{+}21]. Knights [SGC^{+}16]. Know [ABD^{+}97, hH98]. Knowing [hH98].

L2 [Bur97]. L3 [RMC04]. Labs [MD20]. Laguna [Buc97]. Lagunita [WBC^{+}96]. Lake [AFK^{+}19]. Landing [SGC^{+}16]. Landscape [Eec15]. Large [KO05, KKSS99, TKS^{+}99, Yea96]. Large-Scale [KO05]. Larger [RMC04]. Latency [Joh98, Yea96]. latency-hiding [Yea96]. Law [RSS98]. Learning [DPY18, Lie23, MRC^{+}20]. Level [FZW^{+}12, KSI^{+}96]. Lever [Mey06]. Lighting [Tre97]. Lightning [Kre98]. like [Gar95]. Limit [Mei97]. Line [FH00, DGR99]. Liquid [BD99]. Liquid-Crystal-on-Silicon [BD99]. Living [vdWAB^{+}06]. LIW [PSW91]. Llano [BFS12]. LongRun [Fle00]. Look [Lie23]. Looking [Joy96]. Lookup [O’C00a]. Lookups [O’C00b]. Low [ACD^{+}00, BCD^{+}11, CCA^{+}19, Kru00, Luc99, NIJ^{+}03, Plu00a, Plu00b, RC13, SLR^{+}99, WAA^{+}20, Yeh06, Ano93b, Ano94a, Ano94c]. Low-cost [SLR^{+}99, Ano93b]. Low-Power [BCD^{+}11, Kru00, NIJ^{+}03, WAA^{+}20, Yeh06]. Low-Voltage [CCA^{+}19]. LSI [MKN^{+}98].

M3 [RBGZ19]. M32Rx [Shi98]. M32Rx/D [Shi98]. M7 [AJK^{+}15]. Machine [DPY18, MRC^{+}20, Ros99, WP97, TO96b]. Machine-Learning [DPY18]. main [KSI^{+}96]. Mainframe [SBJ13, Web08]. Mainstream [Tre98]. MAJC [Tre99]. Making [SL00]. Management [Fle00, FAWR^{+}11, RNA^{+}12]. Manticore [ZSB21]. Many [FZW^{+}12]. Many-Core [FZW^{+}12]. Manycore [MFN^{+}17]. MAP [Kec97]. MAP1000A [O’D99, BLO00]. Marvell [SSR21]. MasPar [Ano93d]. Massively [BJ06, McC99]. Massively-Parallel [BJ06]. MATRIX [DHM97]. MAX [Lee97]. MAX-2 [Lee97]. Mbps [FH99]. MC68060 [CEM^{+}95]. Mechanisms [DSK^{+}92]. Media [Bat97, CRTI00, DDC^{+}98, Gan98, Kal96, KDK^{+}01, KGM^{+}00, NMP^{+}96, Rat98, SLR^{+}99, SBS^{+}06, vES98, GTB99]. Media-enhanced [KGM^{+}00]. mediaDSP [SP09]. Mediaprocessing [Dub97]. Mediaprocessor [BLO00, Luc99, O’D99, SRD96, THT^{+}04]. Member [McM95]. Memorial [IEE98, IEE13]. Memory [AHM^{+}00, CKD^{+}10, DD05, EGL^{+}90a, FSP06, Joh98, KKK^{+}99, LATSK06, Naa95, NABR95, SB23, TSW^{+}01, WBC^{+}95]. MEMS [TP10]. Merced [War97]. Message [DSK^{+}92, JW98]. Message-Driven [DSK^{+}92]. Metaflow [PSS^{+}91]. Methodologies [DXT^{+}18]. Methods [Gar95]. MHz [Ano93a, Bur97, Gan98, Kes98, LL98, Nt97, SHMS95]. Micro [Mat97]. Microarchitecture [DKyL^{+}17, Kag96, KM03, MS03, RNA^{+}12, TSW^{+}23, Pap96]. Microcontroller [Shi98]. Microlithography [Won03]. Micromachining [Bos96]. micron [BB96]. Microprocessor [ABSS95, ANUN97, ANUN98, ABIK95, BB96, Bur97, Cho98, Chr95, CES^{+}11, Eec15, Joy96, Kes98, KS90, NTK^{+}97, OWJF98, OS08, PSW91, Phi98, RSS^{+}08, SBJ13, Sla97, SAC^{+}98, TKM^{+}02, Web08, Ano93a, Ano94b, ERPR95, JA96, Yea96]. Microprocessors [Eec17, Gar95, LWC^{+}16, LCP^{+}11, Lee95]. Microsoft [Wha98]. milestones [Ano00]. Millenium [Kut99]. Millennium [Tre99]. MIMD [BJ06]. Mining [BH15]. MiniRISC [CC95]. MIPS [CC95, MWV92, Yea96]. MIPS64 [Kru00]. Mitigating [Joh98]. ML [Joh20, MKN^{+}98]. MLPerf [MRC^{+}20]. MMX [Kag96, Mod97, Wei96]. Mobile [BBC^{+}15, CAV^{+}14, Dav02, Dit00, GHSV^{+}11, KKO06, TSI06]. Modular [Oru94]. Monitor [Ros99]. Monolithic [ACD^{+}00]. Montecito [MB05]. Mops [PSW91]. Morello [GBW^{+}23]. Morning [Prz97, Sha96]. Mote [WAP00]. Motion [KSIA95]. Motorola [Als90]. Mountain [FD04]. move [KKO06]. MP [Ano93d, MKN^{+}98, MKN^{+}98]. MP-2 [Ano93d]. Mpact [Kal96]. MPC105 [WBC^{+}95]. MPEG [KSI^{+}96, Nga95, Ste95]. MPEG-1 [Ste95]. MPEG-2 [KSI^{+}96, Nga95]. MPEG2 [KSIA95, MKN^{+}98]. MPLS [ACD^{+}00]. MSP [NMP^{+}96]. Multi [Hes07, MKN^{+}98, MD06, SBKK99, Wei00]. Multi-chip [MKN^{+}98]. Multi-Core [Hes07]. Multi-Standard [MD06]. Multi-Terabit [Wei00]. Multi-Threaded [SBKK99]. Multicomputer [DSK^{+}92]. Multicomputers [PSW91]. Multicore [HWG^{+}09, LATSK06, SP09]. Multimedia [ASK97, ANUN97, ANUN98, Buc97, CAV^{+}14, Dja96, HYYS96, KR96, KBN16, Lee97, Mah96, SANK98, Tre95, vES98, Ano95, KKO06, Lee95, TO96a]. Multiple [PAA^{+}06]. Multiplexed [Jam90]. Multiprocessing [ABG^{+}16, KO05, MD06]. Multiprocessor [KMAC03, NIJ^{+}03, SC91]. Multiprocessors [AAW^{+}96]. Multitenant [MFN^{+}17]. Multithreaded [BBSG11, CSM^{+}21, KST04, KML04, KAO05]. MXi [Jay98]. MxP [CRTI00]. MXT [AHM^{+}00, AHM^{+}00, TSW^{+}01].

N1 [PSB^{+}20]. Named

[DKyL^{+}17, RNA^{+}12]. Names [Vin07]. Native [Gar95]. Near [DFG^{+}13, Khu96]. Near-Threshold [DFG^{+}13]. Neon [MMG^{+}98]. Neoverse [PSB^{+}20]. Netburst [KM03]. Network [Bro00, FH00, GHY^{+}17, Hed00, KML04, NH00, O’C00a, SF18, BWBJ11, SL00]. Network-Facing [KML04]. Networking [CEF^{+}99, FH99]. Networks [ACD^{+}00, CRTI00]. Neural [GHY^{+}17]. New-Generation [MYK^{+}10]. News [Ano00, Ano03, Mat97]. Nexperia [KKO06, Pro06]. Next [AJK^{+}15, AFK^{+}19, BT24, DAV06, ESG^{+}05, KSSF10, KKK^{+}99, PSB^{+}20, SSR21, Vit00, Web08, YHT^{+}15, IDTS00]. Next-Gen [PSB^{+}20]. Next-Generation [AJK^{+}15, BT24, ESG^{+}05, KSSF10, SSR21, Web08, YHT^{+}15]. Niagara [KAO05]. Nintendo64 [Hay97]. nm [ABG^{+}16, CCA^{+}19, FME18, KBN16, Man09, RDJ^{+}13]. Node [DSK^{+}92]. Non [AEJ^{+}00]. Non-Blocking [AEJ^{+}00]. NorthBridge [CH06, AHM^{+}00, KKK^{+}99, OS08]. note [Joh90]. Notebook [Rei96]. Nothing [hH98]. Novel [vES98]. NPU [SL00]. NS486 [Nem95]. NT [Rub97]. Number [Bir98]. NVIDIA [Bur20, CGG^{+}21, Cho23, LNOM08, BBTV15]. NVLink [FD17].

O [Ber98, WAA^{+}20]. Objective [BB96]. Octocore [MYK^{+}10]. off [Wei96]. offs [Pap96, SMHB91]. On-Chip [Bur97, PAGC^{+}97, TO96a]. On-line [DGR99]. One [Sel18, Cra90, SO14]. Open [CCA^{+}19, DXT^{+}18]. Open-Source [CCA^{+}19, DXT^{+}18]. Opening [BWJ98, BBL99, SBS97]. Operating [Fle00, RDJ^{+}13]. Operation [CCA^{+}19]. Opportunities [Rab06]. Opteron [CH06, CKD^{+}10, KMAC03, KO05]. Optical [WAA^{+}20]. Optimization [Kid14, Kni99b, Plu00b]. Optimized [CAV^{+}14]. Optimizing [Pap95, SL00]. Oracle [AJK^{+}15, FJL^{+}13]. Orca [Ari96]. Order [BVD^{+}99, CCA^{+}19, Kes98, Kum96]. Oscillators [TP10]. Other [Alt14, Hun97]. Out-of-Order [BVD^{+}99, CCA^{+}19, Kes98, Kum96]. Outsider [Ano18]. Overview [Bro00, Buc97]. Owns [vdWAB^{+}06].

P55C [Kag96]. P6 [Pap95]. PA [Kum96, Joh98, Lee97, Naa95]. PA-8000 [Kum96, Naa95]. PA-8500 [Joh98]. PA-RISC [Lee97]. Package [WAA^{+}20]. Packet [ACD^{+}00]. Palo [IEE98]. Panel [Bas00, GTB99, JSR^{+}99, War97, WBC^{+}96, Wha98, vdWAB^{+}06]. Papers [RY21]. Parallel [BSP^{+}17, BJ06, CFK^{+}10, McC99, Bar97]. Parallelism [DD05, FZW^{+}12]. Parallelizing [ADG^{+}96, TGK^{+}96, AAW^{+}96]. Part [EGL^{+}90a, Ste90a, Ste90b, EGL^{+}90b]. partially [Joh90]. Parts [Plu00a]. Pascal [FD17]. Pathways [Ano18]. PC [KR96, Tre97, Tre98]. PCI [SRD96, Luc99, SLR^{+}99, WBC^{+}95]. PCs [Ros99]. PE [Ano93d]. Pensando [GM21]. Pentium [Ano94a, Pap96]. PentiumAE [Mod97]. Performance [ASN^{+}99, Bat97, BBSG11, Cha96, CCE^{+}09, CGF18, CGG^{+}21, Cho23, CM00, Cum04, DD05, EG95, FD17, Gol00, GBW^{+}23, HBG^{+}97, Jay98, Khu96, Kru00, Kum96, LCP^{+}11, MRC^{+}20, Mod97, Naa95, Plu00a, Plu00b, Pot97, RC13, Wei96, Yeh06, YHT^{+}15, Ano94a, IEE93, IEE94, OB91, Pap96, TO96a]. Performance/Low [Plu00b]. PERMEDIA [Tre98, Tre96]. Personal [Ano03]. Perspective [AAW^{+}96]. Petaflop [SB23]. Petascale [MYK^{+}10]. Phi [SGC^{+}16]. Philips [KKO06]. Phone [FH00]. Phoneline [CEF^{+}99, FH99]. PicoJava [TO96b]. Pinnacle [TSW^{+}01]. Pioneer [Alt11]. Pipeline [Bat97, FCD^{+}99, Pap95]. Piton [MFN^{+}17]. PivotPoint [Cum04]. Pixel [PAGC^{+}97]. Planet [BDH03]. Platform [ABG^{+}16, KTP^{+}99, Man09, NIJ^{+}03, PSB^{+}20, SP09, GBW^{+}23, Ros99]. PNX4103 [KKO06]. PNX8535 [Pro06]. Point [BSC^{+}90, DKB^{+}90, ZSB21]. Policy [IDTS00]. Port [Bro00]. Portable [LS98, Sla97, THT^{+}04]. Potential [ML21]. Power [ACD^{+}00, BWBJ11, BCD^{+}11, BvdGM^{+}15, DD05, Fle00, Hay97, Hed00, Kid14, Kru00, MM96, NIJ^{+}03, Plu00a, Plu00b, RC13, RNA^{+}12, WAA^{+}20, Yeh06, Ano94a, Ano94c]. Power-Efficient [BvdGM^{+}15]. Power-Management [RNA^{+}12]. POWER10 [STSM21]. POWER4 [FCD^{+}99, MBB^{+}99, Pet00, BTR02]. Power5 [KST04]. Power6 [RSS^{+}08]. Power7 [FAWR^{+}11, KSSF10]. Power9 [STKS17]. Powerful [KTP^{+}99]. PowerPC [Ano94b, Bal95, Bur97, SBKK99]. Preliminary [WAP00]. Preview [DRM^{+}23]. PRISM [SL00]. Pro [Pap96]. Proceedings [Rei96]. Process [Kid14, NPY^{+}21]. Processing [BBC^{+}15, BCF^{+}14, BvdGM^{+}15, DSK^{+}92, DDC^{+}98, KDR^{+}00, KDK^{+}01, KBN16, NMP^{+}96, SHMS95, She99a, She99b, SBS^{+}06, TUHwH99]. Processor [AJK^{+}15, AFK^{+}19, ASK97, Awa95, BH15, Bat97, BBTV15, BSP^{+}17, Bro00, BWBJ11, BCD^{+}11, BvdGM^{+}15, CCA^{+}19, CRTI00, CCE^{+}09, CC95, CKD^{+}10, DSK^{+}92, EGL^{+}90b, EGL^{+}90a, Eer97, EG95, Faa98, FZW^{+}12, FJL^{+}13, Gan98, Gon00, HMB^{+}14, Hay97, Hed00, HYYS96, HWG^{+}09, Jay98, Joh98, Kal96, KST04, KSSF10, KML04, KMAC03, KKO06, KAO05, KGM^{+}00, Kum96, KKSS99, LL98, Mah96, MYK^{+}10, MFN^{+}17, MS03, MB05, MD20, MWV92, Mod97, Nem95, O’C00b, O’C00a, OG01, OKN^{+}00, OW01, Oru94, Pro06, Qua00, Rat98, RDJ^{+}13, RBGZ19, RMC04, STKS17, SHMS95, Sam00a, San96, Sav98, Seg99, Sha00a, Sha00b, She95, SLR^{+}99, SKW^{+}23, STSM21, SVC01, SBKK99, SSB20, SSR21, SANK98, TKS^{+}99, Tre95, Tre97, Yeh06, YHT^{+}15, Ano93b, Ano94a, Chr96, CVS^{+}00, TO96a, GHSV^{+}11, Plu00a]. Processors [BYM^{+}06, CM00, Gon99, Lee97, NH00, SP09, SL00, WMSH09, vES98, Ano95, PAA^{+}06]. Product [SGC^{+}16, Pap96]. Profile [CHH^{+}98, KSI^{+}96]. Profile-Directed [CHH^{+}98]. Profiling [ABD^{+}97, BVD^{+}99]. Program [DGR99, JW98]. Programmability [CGF18]. Programmable [ABG^{+}16, FME18, Rat98, SP09, SBS^{+}06]. Programming [LATSK06]. Programs [TKM^{+}02, AAW^{+}96]. Project [CFO^{+}18]. Prominence [Ano18]. Property [RSS98]. Protocol [NABR95]. Provider [ACD^{+}00]. Pseudo [SHMS95]. Public [Bir98, ESG^{+}05]. Public-Key [ESG^{+}05]. Purpose [ESG^{+}05, MD20, TKM^{+}02]. Purpose-Built [MD20]. PVP [Bar97]. PWRficient [Yeh06]. Pyramid3D [Eer97].

Q [HOF^{+}12]. quality [KKO06]. Quantum [SVC01, CVS^{+}00]. Qubit [SVC01, CVS^{+}00]. Quiet [Cha96].

R10000 [ABSS95, Yea96]. R3D [Pot97]. R3D/100 [Pot97]. R4000 [MWV92]. Rainbows [Vin07]. RAM [FSP06]. Rambus [ASK97, Buc97]. RAMP [PAA^{+}06]. Random [Bir98]. Range [KSIA95, RDJ^{+}13]. ratios [AAW^{+}96]. Raw [TKM^{+}02]. Ray [Pfi99]. Ray-Casting [Pfi99]. Real [CFO^{+}18, Eer97, KSIA95, Pfi99, Yea96]. Real-Time [KSIA95, Pfi99, Eer97]. real-world [Yea96]. Reality [Hay97]. Really [Ano03]. Reconfigurable [DHM97, YYA06]. record [IEE95, IEE96, IEE98]. recorded [AAW^{+}96]. Redstone [DGR99]. Reducing [RC13]. Reduction [Kid14]. Refining [Pap96]. Reliability [BTR02, Qua00]. Reliable [CD95]. rely [Ano95]. Remarks [BWJ98, BBL99, SBS97]. Remembering [Alt11]. Renaissance [KB20]. Research [KZ13, PAA^{+}06]. Resilient [CCA^{+}19]. Revisited [Ste11]. Revolution [Cha96, DPY18, Sam00b]. RF [FME18]. RFSoC [FME18]. Riches [Eec16]. RISC [Gan98, Kum96, OB91, Ano94b, ASK97, ANUN97, ANUN98, Ari96, DXT^{+}18, Gar95, HWG^{+}09, Joh90, KKSS99, Lee97, LWC^{+}16, SHMS95, SB23, SBKK99, SANK98, TKS^{+}99, ZSB21]. RISC-like [Gar95]. RISC-V [DXT^{+}18, LWC^{+}16, SB23, ZSB21]. RISCs [Pen90]. Rock [Ano03, CCE^{+}09]. Room [vdWAB^{+}06]. Route [WBC^{+}96]. Routing [Wei00]. RTX [Bur20]. Run [Ano03, Yea96]. Ryzen [SKP24].

S [SAC^{+}98, OKN^{+}00, Oru94]. S/390 [SAC^{+}98]. S3 [Ste95]. SA [Gan98, Sla97]. SA-1100 [Sla97]. SA-1500 [Gan98]. Samsung [RBGZ19]. Sandy [RNA^{+}12]. Save [Ano03]. SB [Kru00]. SB-1 [Kru00]. Scalable [AFK^{+}19, BCC^{+}02, HWG^{+}09, LL98, NTK^{+}97, SBB^{+}17]. Scalable-Delay-Insensitive [NTK^{+}97]. Scale [Bar21, CFO^{+}18, KO05, KKSS99, TSW^{+}23, TKS^{+}99, VJFG17]. Scale-Out [VJFG17]. Scales [FJL^{+}13]. Scaling [Cho23, FD04, MD20]. Scientific [KKSS99, TKS^{+}99]. Scorpio [Sel18]. Scylla [NABR95]. Search [BDH03]. Second [Phi98, She95, SGC^{+}16]. Second-Generation [SGC^{+}16]. Secrets [Sha96]. Security [GBW^{+}23, Sti19]. Selected [KZ13]. Self [TSV^{+}20, Wei00]. Self-Driving [TSV^{+}20]. Semiconductor [Plu00a, Ano00]. Sensor [Bos96, PAGC^{+}97, SO14]. Sequential [BD99]. Serial [Wei00]. Series [GA21, SKP24]. Server [Ano03, BT24, KSSF10, PKB^{+}15, SGG^{+}12, SSR21]. Server-on-a-Chip [SGG^{+}12]. Servers [Ari96, BCC^{+}02, KMAC03, VJFG17]. ServerSet [AHM^{+}00]. ServerWorks [AHM^{+}00]. Service [ACD^{+}00]. Services [GM21]. Serving [CFO^{+}18]. Session [War97, WBC^{+}96, Wha98]. Set [Awa95, Bre10, FH00, Mah96, Pot97]. SH4 [ANUN97, ANUN98]. Shared [NABR95]. Signal [KDR^{+}00, She99a, She99b]. Silicon [BD99, FD04, GHSV^{+}11, O’C00a, TP10, Tre96, WBC^{+}96]. SIMD [McC99, Phi98, Tha99]. Simple [BYM^{+}06, KSI^{+}96]. simulations [Pap96]. Single [FM95, hH98, MKN^{+}98, NIJ^{+}03, Oru94, Rat98, SC91, Shi98, Ste95]. Single-board [MKN^{+}98]. Single-Chip [NIJ^{+}03, SC91, MKN^{+}98]. SiRFstar [KTP^{+}99]. Sitera [SL00]. Sixth [Mah96]. Sixth-Generation [Mah96]. Skullduggery [Ste11]. Sky [Mei97]. Skylake [DKyL^{+}17]. Small [TSI06]. Smaller [Bal95]. Smart [PAGC^{+}97, SF18, WAP00]. SmartNIC [DRM^{+}23]. SMP [KKSS99, TKS^{+}99]. SOC [Plu00a, CM00, PSB^{+}20]. Sockets [FJL^{+}13]. Software [AAW^{+}96, BJ06, GHY^{+}17, Joh90, Kal96, Lie23, Sha96, SKW^{+}23, TKM^{+}02, WBC^{+}96, Ano95, VBC^{+}21]. Software-Hardware [GHY^{+}17]. Soggy [Joh90]. Solution [TSV^{+}20]. Solutions [Won03]. Sort [ASN^{+}99]. Sorting [Prz97]. Source [CCA^{+}19, DXT^{+}18]. Sparc [FJL^{+}13, CCE^{+}09, Nt97, She95, AJK^{+}15, BSC^{+}90, BAC^{+}90, DKB^{+}90, KAO05, SGG^{+}12]. SPARC64 [EG95, She95, MYK^{+}10, YHT^{+}15]. Speaker [Pap98]. SPECfp [AAW^{+}96]. Special [DA23, LV24, SS22]. Specializer [DGR99]. Spectrum [DDC^{+}98]. Speculative [EG95]. Speed [BWBJ11, Kre98, Luc99, O’C00a, TP10, SLR^{+}99]. speedAI240 [SB23]. Sphere [Sti19]. Spillovers [Gre11]. springs [Joh90]. SR8000 [KKSS99, TKS^{+}99]. Stacked [DFG^{+}13]. Standard [MRC^{+}20, MD20, MD06, NMP^{+}96, Par98, Ano00]. Standardization [Ste11]. Stanford [IEE93, IEE94, IEE95, IEE96, IEE97, IEE98, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, IEE13, Rei96, Ano94c, CFK^{+}10, HHS^{+}99, HHS^{+}00]. Start [Bas00]. Start-ups [Bas00]. Steroids [HBG^{+}97]. Storage [CD95, Dav02]. StrataSwitch [EM99]. Streaming [Faa98, Tha99]. Streams [KDR^{+}00, KDK^{+}01]. StrongARM [Sla97, San96]. StrongARMing [LS98]. Stuff [BS98]. Substrate [VBC^{+}21]. Subsystem [CKD^{+}10, KKK^{+}99]. success [Joh90]. Sugar [Ano03]. Suite [MRC^{+}20]. Sun [HBG^{+}97]. Superscalar [ABSS95, ASK97, CEM^{+}95, CC95, ERPR95, EG95, LL98, McM95, SHMS95, SANK98, Tre95, Yea96]. Support [BVD^{+}99, NABR95, Tre95, Ano95]. supporting [TO96a]. Surface [Bos96]. Switch [AEJ^{+}00, Cum04, Edd02, EM99, Wei00]. Switches [IDTS00]. Switching [KSI^{+}96]. Symmetric [Bir98, KO05]. Symposium [HW91, IEE93, IEE94, IEE95, IEE96, Ste90a, Ste90b]. Synchronous [FCD^{+}99, Wei00]. Synthesis [KIS^{+}99, KIS^{+}00]. Synthesizable [Bur99, Seg99]. System [ABG^{+}16, BTR02, BCF^{+}14, BWBJ11, CES^{+}11, DRM^{+}23, DFG^{+}13, Fos98, Gar00, Gon99, GBW^{+}23, GA21, KTP^{+}99, NCT^{+}98, SO14, SBS^{+}06, WMSH09, Ano93c, Joh90, Ari96, OB91]. System-On-Chip [SBS^{+}06, ABG^{+}16, Gar00]. System/6000 [Ari96, OB91]. Systems [Cum04, DKB^{+}90, Joh19, KO05, LL98, MD20, OKN^{+}00, OW01, TP10, TGK^{+}96].

T [CSM^{+}21, FZW^{+}12]. T4 [SGG^{+}12]. T5 [FJL^{+}13]. Table [O’C00b, O’C00a]. Take [Joh90]. Task [BSP^{+}17]. Task-Parallel [BSP^{+}17]. Technical [Mal97]. Techniques [Joh98, Kni99b, Yea96]. Techniques/Using [Kni99b]. Technology [AHM^{+}00, Bos96, DRM^{+}23, Dav02, DDC^{+}98, Fan99a, Fan99b, HMR96, Kag96, KM03, Mat11, Mey06, Mod97, OWJF98, Phi98, WAA^{+}20, Wei96]. TechPress [ES99, GTB99]. Television [Pro06]. Tensor [CGG^{+}21]. Terabit [Wei00]. Teraflops/W [SB23]. TeraOPS [BJ06]. TeraPHY [WAA^{+}20]. Tesla [LNOM08, TSV^{+}20, TSW^{+}23]. Testchip [MBB^{+}99]. Their [Won03]. Them [Alt13]. Themes [Alt14]. Third [SBJ13, Tre98]. Third-Generation [SBJ13]. Thread [FZW^{+}12, MB05]. Thread-Level [FZW^{+}12]. Threaded [SGG^{+}12, SBKK99]. Threads [LATSK06]. Threshold [DFG^{+}13]. ThunderX3 [SSR21]. Tiered [DXT^{+}18]. Time [CFO^{+}18, KSIA95, Pfi99, Eer97]. TITAC [NTK^{+}97]. TITAC-2 [NTK^{+}97]. TM [SRD96, SLR^{+}99]. TM-1 [SRD96]. TM-1300 [SLR^{+}99]. TM1300 [Luc99]. TMS320 [Gol00]. TMS320C6xxx [Tru97]. TMS390C602A [DKB^{+}90]. Tolerant [RSS^{+}08]. totally [Ano93d]. Touchstone [Dja96, KR96]. Tough [Ano95]. TPUv2 [NPY^{+}21]. TPUv3 [NPY^{+}21]. Trade [Wei96, Pap96, SMHB91]. Trade-off [Wei96]. trade-offs [Pap96, SMHB91]. Traditional [LCP^{+}11]. Training [MD20, NPY^{+}21]. Transactional [LATSK06]. Transceiver [GDES08]. Transistor [RC13]. Transistors [Kid14]. Translator [CHH^{+}98]. Transmeta [Dit00]. Transparent [Rub97]. Trends [Won03]. Trimedia [Luc99, SRD96]. True [Vin07]. Tuning [Pap96]. Turing [Bur20]. Turns [KvdW09]. Tutorial [Cas00, Dub97, KFL99, NH00, Pra96, Prz97, Sha96]. Two [KSIA95, KSI^{+}96, Par98]. Two-Chip [KSIA95, KSI^{+}96].

Ubiquitous [CFK^{+}10]. Ultra [FSP06, FD17, TSI06, Ano94c]. Ultra-dense [FSP06]. Ultra-Performance [FD17]. Ultraefficient [ZSB21]. UltraSPARC [ADG^{+}96, LL98, NCT^{+}98, TGK^{+}96, Tre95, Nt97, TO96a]. UltraSPARC-I [Tre95]. UltraSPARC-IIi [NCT^{+}98]. Uncertain [WD03]. Uncompressed [GDES08]. Unified [LNOM08]. Uniprocessors [Pra96]. Unit [BBC^{+}15, BCF^{+}14, KBN16, KIS^{+}00]. Units [KIS^{+}99]. University [IEE93, IEE94, IEE95, IEE96, IEE97, IEE98, IEE99, IEE00, IEE01, IEE02, IEE03, IEE04, IEE05, IEE06, IEE07, IEE08, IEE09, IEE10, IEE11, Rei96]. ups [Bas00]. Using [KDR^{+}00, Kid14, Kni99b, MD20, O’C00a, O’C00b, YYA06]. Utility [Rub97].

V [IEE93, DXT^{+}18, GDES08, LWC^{+}16, SB23, ZSB21]. V830R [ASK97, SANK98]. V830R/AV [ASK97, SANK98]. V9 [Nt97]. Validating [GBW^{+}23]. Value [Ari96]. Variability [RC13]. Vector [ABIK95, CSM^{+}21, Faa98, KGM^{+}00, KIS^{+}99, KIS^{+}00, SHMS95, SBB^{+}17, KGM^{+}00]. VelaTX [LD98]. VelociTI [Tru97]. Verification [EGL^{+}90a]. versus [Bas00]. vg500 [Pfi99]. VI [IEE94, AS95]. Video [Bar97, CM00, FM95, KSIA95, ML21, MKN^{+}98, MD06, Nga95, SC91, SP09, Ste95, Vis99, KSI^{+}96]. VII [IEE95]. VIII [IEE96, Rei96]. VIIIfx [MYK^{+}10]. ViRGE [Khu96]. Virtex5 [DAV06]. Virtual [Ros99, TO96b, WP97, Ros99]. Vision [BBC^{+}15]. Visiting [Mat97]. Vitesse [SL00]. VLIW [BLO00, Gar95, HYYS96, Luc99, O’D99, SRD96, SLR^{+}99, vES98]. VLSI [Dja96, Nga95]. VMware [Ros99]. Voice [WMSH09]. VoIP [CRTI00]. Volta [CGF18]. Voltage [CCA^{+}19, RDJ^{+}13]. vs [Gar95]. VX [Khu96].

W [SB23]. Wabi [HMR96]. Warehouse [Bar21]. Warehouse-Scale [Bar21]. Wars [Jam90]. Wave [FCD^{+}99, Ano95]. Wave-Pipeline [FCD^{+}99]. Way [Bas00, KAO05, LL98]. Web [BDH03]. Welcome [BWJ98, BBL99, SBS97]. Well [hH98]. Were [War97]. Where [ABD^{+}97]. While [hH98]. whilst [KKO06]. Who [vdWAB^{+}06]. Wide [KSIA95, O’C00b, O’C00a, RDJ^{+}13]. Wide-Voltage-Operating [RDJ^{+}13]. Wiggins [DGR99]. Wiggins/Redstone [DGR99]. will [Ano95]. Win32 [Rub97]. Wire [BWBJ11]. Wire-Speed [BWBJ11]. Wireless [GDES08, Plu00b, Rab06, Vit00]. Within [ML21]. without [Chr96]. Workloads [KML04]. Workstation [Khu96, MMG^{+}98, Ano93c]. World [Sam99, Yea96]. WTL3170 [BSC^{+}90]. WTL3170/3171 [BSC^{+}90].

X [Sel18, GA21]. x86 [BCD^{+}11, Hes07, HWG^{+}09, Jay98, Mah96, Chr96, Fle00, Gar95, Rub97]. Xbox [GA21, SO14, Sel18]. xDSL [She99a]. Xeon [SGC^{+}16, AFK^{+}19]. XIfx [YHT^{+}15]. Xtensa [Gon00]. XXX [Pap98].

yield [AAW^{+}96].

Z [FSP06]. Z-RAM [FSP06]. z10 [Web08]. zEC12 [SBJ13]. Zen [BT24, SSB20]. zEnterprise [CES^{+}11].

References

Amarisinghe:1995:HCF

[AAFH95] S. P. Amarisinghe, J. A. M.

Anderson, R. S. French, and M. W. Hall. Hot compilers for future hot chips. In IEEE [IEE95], pages 167–178. ISBN ???? LCCN ????

Amarasinghe:1996:MSP

[AAW^{+}96] Saman P. Amarasinghe, Jennifer M. Anderson, Christopher S. Wilson, Shih-Wei Liao, Brian R. Murphy, Robert S. French, Monica S. Lam, and Mary W. Hall. Multiprocessors from a software perspective — automatically parallelizing benchmark programs to yield the highest SPECfp ratios recorded. IEEE Micro, 16(3): 52–61, May/June 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Anderson:1997:CPD

[ABD^{+}97] Jennifer Anderson, Lance Berc, Jeff Dean, Sanjay Ghemawat, Monika Henzinger, Shun-Tak Leung, Dick Sites, Mitch Lichtenberg, Mark Vandevoorde, Carl Waldspurger, and Bill Weihl. Continuous profiling: (it’s 10:43; do you know where your cycles are?). In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Ahmad:2016:NMS

[ABG^{+}16] Sagheer Ahmad, Vamsi Boppana, Ilya Ganusov, Vinod Kathail, Vidya Rajagopalan, and Ralph Wittig. A 16-nm multiprocessing system-on-chip field-programmable gate array platform. IEEE Micro, 36(2): 48–62, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2016/02/mmi2016020048-abs.
html.

Asanovic:1995:VM

[ABIK95] K. Asanovic, J. Beck, B. Irissou, and B. E. D. Kingsbury. The TO vector microprocessor. In IEEE [IEE95], pages 187–196. ISBN ???? LCCN ????

Ahi:1995:RSM

[ABSS95] A. Ahi, A. Bomdica, G. Shippen, and H. Sucar. R10000 superscalar microprocessor. In IEEE [IEE95], pages 227–238. ISBN ???? LCCN ???? URL

ftp://ftp.sgi.com/sgi/doc/
R10000/hotchips/hotchips.
ps;   ftp://ftp.sgi.com/sgi/
doc/R10000/hotchips/hotchips_
talk.ps.

Asami:2000:LPM

[ACD^{+}00] T. Asami, E. Chao, N. Dagli, J. Dickinson, J. Fiorenza, S. Fallow, P. Gopi, O. Hassen, J. Hamada, N. Hudson, R. Krishnan, H. Luu, C. Nabangxang, K. Peng, D. Peeters, P. Wang, A. Yoshida, and J. Zoll. A low power monolithic IP/MPLS packet forwarding engine for service provider networks. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http:/

/www.hotchips.org/index12.
html.

Aoki:1996:PCU

[ADG^{+}96] Chris Aoki, Peter Damron, Kurt Goebel, Vinod Grover, Xiangyun Kong, Michael Lai, Krishna Subramanian, Partha Tirumalai, and Jian-Zhong Wang. A parallelizing compiler for UltraSPARC. In IEEE [IEE96], page ?? ISBN ???? LCCN ????

Alowersson:2000:CNB

[AEJ^{+}00] Jonas Alowersson, Anders Edman, Henrik O. Johansson, Tomas Johansson, Anders Lloyd, Bertil Roslund, Lars-Olof Svensson, Patrik Sundström, Peter Tufvesson, Kenny Ranerup, Per Andersson, and Christer Svensson. A CMOS non-blocking 16 × 16 Gigabit Ethernet switch chip. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Arafa:2019:CLN

[AFK^{+}19] M. Arafa, B. Fahim, S. Kottapalli, A. Kumar, L. P. Looi, S. Mandava, A. Rudoff, I. M. Steiner, B. Valentine, G. Vedaraman, and S. Vora. Cascade Lake: Next generation Intel Xeon scalable processor. IEEE Micro, 39(2):29– 36, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arramreddy:2000:IMM

[AHM^{+}00] Sujith Arramreddy, David Har, Kwok-Ken Mak, R. Brett Tremaine, and Michael Wazlowski. IBM “MXT” memory compression technology debuts in a ServerWorks Northbridge ServerSet III and MXT technology. In IEEE [IEE00], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/index12.html.

Aingaran:2015:MON

[AJK^{+}15] Kathirgamar Aingaran, Sumti Jairath, Georgios Konstadinidis, Serena Leung, Paul Loewenstein, Curtis McAllister, Stephen Phillips, Zoran Radovic, Ram Sivaramakrishnan, David Smentek, and Thomas Wicki. M7: Oracle’s next-generation Sparc processor. IEEE Micro, 35(2):36– 45, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020036-abs.
html.

Alsup:1990:MFA

[Als90] Mitch Alsup. Motorola’s 88000 family architecture. IEEE Micro, 10(3):48–66, May/June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2011:HCR

[Alt11] Erik R. Altman. Hot Chips and remembering a pioneer. IEEE Micro, 31(2):3, March/ April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2013:HCI

[Alt13] Erik R. Altman. Hot chips and the incomplete job of exploiting them. IEEE Micro, 33(2): 4–5, March/April 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Altman:2014:HCO

[Alt14] Erik R. Altman. Hot chips and other themes. IEEE Micro, 34 (2):2–3, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Amirtharajah:2008:GEI

[AM08] Raj Amirtharajah and John Mashey. Guest Editors’ introduction: Hot Chips 19. IEEE Micro, 28(2):7–9, March/ April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/mi/2008/02/mmi2008020007.
                                        pdf.

Anonymous:1993:MBE

[Ano93a] Anonymous. A 300 MHz 115W 32b bipolar ECL microprocessor. In IEEE [IEE93], page ?? ISBN ???? LCCN ????

Anonymous:1993:DAA

[Ano93b] Anonymous. DECchip21066 — Alpha AXP architecture processor for low-cost applications. In IEEE [IEE93], page ?? ISBN ???? LCCN ????

Anonymous:1993:HIW

[Ano93c] Anonymous. A highly-integrated workstation graphics system. In IEEE [IEE93], page ?? ISBN ???? LCCN ????

Anonymous:1993:MMP

[Ano93d] Anonymous. MasPar MP-2 PE chip: a totally cool hot chip. In IEEE [IEE93], page ?? ISBN ???? LCCN ????

Anonymous:1994:HPL

[Ano94a] Anonymous. A high performance, low power, Pentium processor. In IEEE [IEE94], page ?? ISBN ???? LCCN ????

Anonymous:1994:PRM

[Ano94b] Anonymous. PowerPC 604 RISC microprocessor. In IEEE [IEE94], page ?? ISBN ???? LCCN ????

Anonymous:1994:SUL

[Ano94c] Anonymous. Stanford ultra low power CMOS. In IEEE [IEE94], page ?? ISBN ???? LCCN ????

Anonymous:1995:HCT

[Ano95] Anonymous. Hot chips, tough choices — A new wave of multimedia processors will rely heavily on software support. BYTE Magazine, 20(8):25–??, August 1995. CODEN BYTEDJ. ISSN 0360-5280.

Anonymous:2000:NHI

[Ano00] Anonymous. News: Hot interconnects, hot chips, Infini-Band standard, semiconductor milestones. IEEE Micro, 20 (6):2–3, November/December 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/

mi/books/mi2000/pdf/m6002.
pdf.

Anonymous:2003:NSC

[Ano03] Anonymous. News 2.0: Sugar cubes rock; can a personal server really save your day?; hot chips run cool. ACM Queue: Tomorrow’s Computing Today, 1(7):6, October 2003. CO-DEN AQCUAE. ISSN 1542-7730 (print), 1542-7749 (electronic).

Anonymous:2018:OPP

[Ano18] Anonymous. Outsider pathways to prominence. IEEE Micro, 38(2):81–84, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020081.html.

Arakawa:1997:SRM

[ANUN97] Fumio Arakawa, Osamu Nishii, Kunio Uchiyama, and Norio Nakagawa. SH4 RISC microprocessor for multimedia. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Arakawa:1998:SRM

[ANUN98] Fumio Arakawa, Osamu Nishii, Kunio Uchiyama, and Norio Nakagawa. SH4 RISC multimedia microprocessor. IEEE Micro, 18(2):26–34, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL hf=

0?target=if(eq(query(%27%
3CFNO%3E+cont+m2026%27),0)

,1,ancestor(ARTICLE,query(
%27%3CFNO%3E+cont+m2026%27)
));   http://dlib.computer.
org/dynaweb/mi/mi1998/@ebt-
link;   http://dlib.computer.
org/mi/books/mi1998/pdf/m2026.
pdf;   http://www.computer.
org/micro/mi1998/m2026abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Amirtharajah:2016:HC

[AR16a] Rajeevan Amirtharajah and Behnam Robatmili. Hot Chips 27. IEEE Micro, 36(2):6– 7, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2016/02/mmi2016020006-abs.
html.

Amirtharajah:2016:HCH

[AR16b] Rajeevan Amirtharajah and Behnam Robatmili. Hot chips 27 highlights. IEEE Micro, 36 (2):64–69, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2016/02/mmi2016020064.html.

Arimilli:1996:OCH

[Ari96] Ravi Arimilli. The Orca chip … heart of IBM’s RISC System/ 6000 “Value” servers. In IEEE [IEE96], pages 35–46. ISBN ???? LCCN ????

Alpert:1995:GEI

[AS95] Donald Alpert and Alan Jay Smith. Guest Editors’ introduction: Hot Chips VI. IEEE Micro, 15(2):8–9, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Arai:1997:VAE

[ASK97] Tomohisa Arai, Kazumasa Suzuki, and Ichiro Kuroda. V830R/AV embedded multimedia superscalar RISC processor with Rambus interface. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Azuma:1999:HPS

[ASN^{+}99] Shinsuke Azuma, Takao Sakuma, Takashi Nakano, Takaaki Ando, and Kenji Shirai. Highperformance sort chip. In IEEE [IEE99], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/hotc11_index.
                                        html.

Asanovic:2010:GEI

[AW10] Krste Asanović and Ralph Wittig. Guest Editors’ introduction: Hot Chips 21. IEEE Micro, 30(2):5–6, March/April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Awaga:1995:GPC

[Awa95] M. Awaga. 3D graphics processor chip set. In IEEE [IEE95], pages 121–134. ISBN ???? LCCN ????

Brown:1990:ISE

[BAC^{+}90] Emil W. Brown, Anant Agrawal, Trevor Creary, Michael F. Klein, David Murata, and Joseph Petolino. Implementing Sparc in ECL. IEEE Micro, 10(1):10–22, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Balser:1995:SFC

[Bal95] D. Balser. Smaller, faster, cooler … evolving the PowerPC family. In IEEE [IEE95], pages 217–226. ISBN ???? LCCN ????

Barthel:1997:PPV

[Bar97] Dominique Barthel. PVP: a Parallel Video coProcessor. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Barroso:2021:BHW

[Bar21] L. A. Barroso. A brief history of warehouse-scale computing. IEEE Micro, 41(2):78– 83, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baskett:2000:PSU

[Bas00] Forest Baskett. Panel: Startups versus established companies: The best way to innovate? In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Battle:1997:EHP

[Bat97] Jim Battle. Efficient high performance 3D pipeline implementation on a media processor. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Bouchard:1996:DOM

[BB96] Gregg Bouchard and Pete Bannon. Design objective of the 0.35-micron Alpha 21164 microprocessor. In IEEE [IEE96], pages 21–34. ISBN ???? LCCN ????

URL ftp://www.hotchips.
org/pub/hotc7to11cd/hc96/
hc8_pdf/1.2.pdf.

Baum:2012:HC

[BB12] Allen Baum and Bevan Bass. Hot Chips 23. IEEE Micro, 32 (2):6–7, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barry:2015:AVP

[BBC^{+}15] Brendan Barry, Cormac Brick, Fergal Connor, David Donohoe, David Moloney, Richard Richmond, Martin O’Riordan, and Vasile Toma. Always-on vision processing unit for mobile applications. IEEE Micro, 35(2):56–66, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://www.

computer.org/csdl/mags/mi/
2015/02/mmi2015020056-abs.
html.

Blasgen:1999:WOR

[BBL99] Michael Blasgen, Forest Baskett, and Monica Lam. Welcome, opening remarks. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Butler:2011:BAM

[BBSG11] Michael Butler, Leslie Barnes, Debjit Das Sarma, and Bob Gelinas. Bulldozer: An approach to multithreaded compute performance. IEEE Micro, 31(2):6–15, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Boggs:2015:DNF

[BBTV15] Darrell Boggs, Gary Brown, Nathan Tuck, and K S Venkatraman. Denver: Nvidia’s first 64-bit ARM processor. IEEE Micro, 35(2):46– 55, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020046-abs.
html.

Briggs:2002:IBB

[BCC^{+}02] Fayé Briggs, Michel Cekleov, Ken Creta, Manoj Khare, Steve Kulick, Akhilesh Kumar, Lily Pao Looi, Chitra Natarajan, Sivakumar Radhakrishnan, and Linda Rankin. Intel 870: A building block for costeffective, scalable servers. IEEE Micro, 22(2):36–47, March/ April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m2036.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m2036abs.
                                        htm.

Burgess:2011:BAL

[BCD^{+}11] Brad Burgess, Brad Cohen, Marvin Denman, Jim Dundas, David Kaplan, and Jeff Rupley. Bobcat: AMD’s low-power x86 processor. IEEE Micro, 31(2): 16–25, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bouvier:2014:KAa

[BCF^{+}14] Dan Bouvier, Brad Cohen, Walter Fry, Sreekanth Godey, and Michael Mantor. Kabini: An AMD accelerated processing unit system on a chip. IEEE Micro, 34(2):22–33, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Bolotski:1999:FSC

[BD99] Michael Bolotski and Jean-Jacques Drolet. A fieldsequential color 1040 by 768 liquid-crystal-on-silicon display. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Barroso:2003:WSP

[BDH03] Luiz André Barroso, Jeffrey Dean, and Urs Hölzle. Web search for a planet: The Google Cluster Architecture. IEEE Micro, 23(2):22–28, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2003/pdf/m2022.
pdf;   http://www.computer.
org/micro/mi2003/m2022abs.
htm.

Berkowitz:1998:FCG

[Ber98] Fred Berkowitz. Fast CPUs are good … but fast I/O is better. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Branover:2012:AFA

[BFS12] Alexander Branover, Denis Foley, and Maurice Steinman. AMD Fusion APU: Llano. IEEE Micro, 32(2):28– 37, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Barkatullah:2015:GCF

[BH15] Javed Barkatullah and Timo Hanke. Goldstrike 1: Coin-Terra’s first-generation cryptocurrency mining processor for Bitcoin. IEEE Micro, 35(2): 68–76, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020068-abs.
html.

Birman:1998:ACH

[Bir98]  Mark Birman.  Accelerat-

ing cryptography in hardware: Public key, random number generation, symmetric key. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Butts:2006:THS

[BJ06] Mike Butts and Anthony Mark Jones. TeraOPS hardware & software: A new massivelyparallel, MIMD computing fabric IC. Web document., 2006.

                                        URL http://www.hotchips.
                                        org/archives/hc18/2_Mon/HC18.
                                        S5/HC18.S5T1.pdf.

Basoglu:2000:MVM

[BLO00] Chris Basoglu, Woobin Lee, and John Setel O’Donnell. The Map1000A VLIW mediaprocessor. IEEE Micro, 20(2):48–59, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2000/pdf/m2048.
                                        pdf;   http://www.computer.
                                        org/micro/mi2000/m2048abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Boser:1996:SMI

[Bos96] Bernhard E. Boser. Surface micromachining: An IC compatible sensor technology. In IEEE [IEE96], pages 241–256. ISBN ???? LCCN ????

Brewer:2010:ISI

[Bre10] Tony M. Brewer. Instruction set innovations for the Convey HC-1 computer. IEEE Micro, 30(2):

70–79, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Brown:2000:CPC

[Bro00] Andy Brown. The C-Port C-5 network processor: An architectural overview. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Baum:1998:GEI

[BS98] Allen J. Baum and Alan Jay Smith. Guest Editors’ introduction: Hot chips—hot stuff. IEEE Micro, 18(2):11– 13, March/April 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL hf=0?target=
if(eq(query(%27%3CFNO%3E+
cont+m2011%27),0),1,ancestor(
ARTICLE,query(%27%3CFNO%3E+
cont+m2011%27)));   http:/
/dlib.computer.org/dynaweb/
mi/mi1998/@ebt-link;   http:
//dlib.computer.org/mi/books/
mi1998/pdf/m2011.pdf.

Birman:1990:DWS

[BSC^{+}90] Mark Birman, Allen Samuels, George Chu, Ting Chuk, Larry Hu, John McLeod, and John Barnes. Developing the WTL3170/3171 Sparc floatingpoint coprocessors. IEEE Micro, 10(1):55–64, January/ February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bohnenstiehl:2017:KFG

[BSP^{+}17] Brent Bohnenstiehl, Aaron Stillmaker, Jon Pimentel, Timothy Andreas, Bin Liu, Anh Tran, Emmanuel Adeagbo, and Bevan Baas. KiloCore: A fine-grained 1,000-processor array for task-parallel applications. IEEE Micro, 37(2):63– 69, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/mi/
                                        2017/02/mmi2017020063-abs.
                                        html.

Bhargava:2024:ANG

[BT24] Ravi Bhargava and Kai Troester. AMD next-generation “Zen 4” core and 4th gen AMD EPYC server CPUs. IEEE Micro, 44 (3):8–17, May/June 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Bossen:2002:PSD

[BTR02] Douglas C. Bossen, Joel M. Tendler, and Kevin Reick. Power4 system design for high reliability. IEEE Micro, 22(2):16–24, March/April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2002/pdf/m2016.
                                        pdf;   http://www.computer.
                                        org/micro/mi2002/m2016abs.
                                        htm.

Buchanan:1997:OLI

[Buc97] Mike Buchanan. Overview of the Laguna II Rambus multimedia accelerator. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Burgess:1997:MPM

[Bur97] Brad Burgess. A 250 MHz 5W PowerPC microprocessor with on-chip L2 cache controller. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Burdass:1999:AAS

[Bur99] A. Burdass. The ARM9TDMI and ARM9ESP synthesizable families. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Burgess:2020:RNT

[Bur20]  J. Burgess.  RTX on the

NVIDIA Turing GPU. IEEE Micro, 40(2):36–44, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Berc:1999:HSO

[BVD^{+}99] Lance Berc, Mark Vandevoorde, J. Dean, S. Ghemawat, S. Leung, M. Litchenberg, M. Vandevoorde, G. Verns, C. Waldspurger, W. Weihl, and J. White. Hardware support for out-of-order instruction profiling on Alpha 21264a. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Burres:2015:IAC

[BvdGM^{+}15] Bradley Burres, Johan van de Groenendaal, Praveen Mosur, Jonathan Robinson, Ian Steiner, Yi-Feng Liu, Sin S. Tan, Erik McShane, Belliappa Kuttanna, and Sridhar Lakshmanamurthy. Intel Atom C2000 processor family: Powerefficient datacenter processing. IEEE Micro, 35(2):26– 34, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020026-abs.
html.

Brown:2011:IPE

[BWBJ11] Jeffrey D. Brown, Sandra Woodward, Brian M. Bass, and Charles L. Johnson. IBM Power Edge of Network processor: A wire-speed system on a chip. IEEE Micro, 31(2): 76–85, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Baum:1998:WOR

[BWJ98] Allen Baum, John Wawrzynek, and Norm Jouppi. Welcome and opening remarks. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Baas:2006:HAA

[BYM^{+}06] Bevan Baas, Zhiyi Yu, Michael Meeuwsen, Omar Sattari, Ryan Apperson, Eric Work, Jeremy Webb, Michael Lai, Daniel Gurman, Chi Chen, Jason Cheung, Dean Truong, and Tinoosh Mohsenin. Hardware and applications of AsAP: An asynchronous array of simple processors. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.S5/HC18.S5T3.
pdf.

Cassiday:2000:IAT

[Cas00] Daniel Cassiday. InfinBand architecture tutorial. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http://

www.hotchips.org/docs/iba_
tutorial_hot_chips.pdf.

Codrescu:2014:HDA

[CAV^{+}14] Lucian Codrescu, Willie Anderson, Suresh Venkumanhanti, Mao Zeng, Erich Plondke, Chris Koob, Ajay Ingle, Charles Tabony, and Rick Maule. Hexagon DSP: An architecture optimized for mobile multimedia and communications. IEEE Micro, 34(2):34–43, March/ April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Cobb:1995:MCS

[CC95] P. Cobb and J. Cesana. The MiniRISC CW4010: A superscalar MIPS processor ASIC core. In IEEE [IEE95], pages 19–30. ISBN ???? LCCN ????

Celio:2019:BOS

[CCA^{+}19] C. Celio, P. Chiu, K. Asanovic, B. Nikolic, and D. Patterson. BROOM: An open-source outof-order processor with resilient low-voltage operation in 28-nm CMOS. IEEE Micro, 39(2): 52–60, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chaudhry:2009:RHP

[CCE^{+}09] Shailender Chaudhry, Robert Cypher, Magnus Ekman, Martin Karlsson, Anders Landin, Sherman Yip, Håkan Zeffer, and Marc Tremblay. Rock: A highperformance SPARC CMT processor. IEEE Micro, 29(2):6– 16, March/April 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cheng:1995:FHR

[CD95] J. M. Cheng and L. M. Duyanovich. Fast and highly reliable IBMLZ1 compression chip and algorithm for storage. In IEEE [IEE95], pages 143–154. ISBN ???? LCCN ????

Corry:1999:EEE

[CEF^{+}99] A. Corry, G. Efland, E. Frank, N. Ferrario, H. Garlapati, R. Hayes, J. Holloway, H. Kuo, J. Laudon, T. Mallory, W. Morton, G. Loyola, N. Nucklos, J. Pattin, H. Ptasinski, K. Peterson, E. Ojard, D. Snow, W. Stafford, T. Robinson, J. Trachewsky, L. Yamano, C. Young, C. Warth, R. Alva, B. Bunch, D. Fifield, N. Castagnoli, M. Dove, M. Kobayashi, R. McCauley, S. Mohapatra, T. Moorti, A. Siddeqee, W. Shieh, and S. Siener. The Epigram EPI41210/EPI41100 16bps home phoneline networking chipset. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Circello:1995:SAM

[CEM^{+}95] Joe Circello, Greg Edgington, Dan McCarthy, James Gay, David Schimke, Steven Sullivan, Richard Duerden, Chris Hinds, Danny Marquette, Lal Sood, Al Crouch, and Daniel Chow. The superscalar architecture of the MC68060. IEEE Micro, 15(2):10–21, March/April 1995. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Curran:2011:ZSM

[CES^{+}11] Brian W. Curran, Lee E. Eisen, Eric M. Schwarz, Pak kin Mak, James Warnock, Patrick J. Meaney, and Michael Fee. The zEnterprise 196 system and microprocessor. IEEE Micro, 31 (2):26–40, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Catanzaro:2010:UPC

[CFK^{+}10] Bryan Catanzaro, Armando Fox, Kurt Keutzer, David Patterson, Bor-Yiing Su, Marc Snir, Kunle Olukotun, Pat Hanrahan, and Hassan Chafi. Ubiquitous parallel computing from Berkeley, Illinois, and Stanford. IEEE Micro, 30(2):41– 55, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chung:2018:SDR

[CFO^{+}18] Eric Chung, Jeremy Fowers, Kalin Ovtcharov, Michael Papamichael, Adrian Caulfield, Todd Massengill, Ming Liu, Daniel Lo, Shlomi Alkalay, Michael Haselman, Maleen Abeydeera, Logan Adams, Hari Angepat, Christian Boehn, Derek Chiou, Oren Firestein, Alessandro Forin, Kang Su Gatlin, Mahdi Ghandi, Stephen Heil, Kyle Holohan, Ahmad El Husseini, Tamas Juhasz, Kara Kagi, Ratna Kovvuri, Sitaram Lanka, Friedel van Megen, Dima Mukhortov, Prerak Patel, Brandon Perez, Amanda Rapsang, Steven Reinhardt, Bita Rouhani, Adam Sapek, Raja Seera, Sangeetha Shekar, Balaji Sridharan, Gabriel Weisz, Lisa Woods, Phillip Yi Xiao, Dan Zhang, Ritchie Zhao, and Doug Burger. Serving DNNs in real time at datacenter scale with Project Brainwave. IEEE Micro, 38(2):8– 20, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020008-abs.
html.

Choquette:2018:VPP

[CGF18] Jack Choquette, Olivier Giroux, and Denis Foley. Volta: Performance and programmability. IEEE Micro, 38(2):42– 52, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020042-abs.
html.

Choquette:2021:NAT

[CGG^{+}21] J. Choquette, W. Gandhi, O. Giroux, N. Stam, and R. Krashinsky. NVIDIA A100 Tensor Core GPU: Performance and innovation. IEEE Micro, 41(2):29–35, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Conway:2006:OCN

[CH06] Pat Conway and Bill Hughes.

The Opteron CMP North-Bridge architecture, now and in the future. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.S2/HC18.S2T2.
pdf.

Chapman:1996:HPC

[Cha96] David Chapman. High performance caches: The quiet revolution. In IEEE [IEE96], pages 95–108. ISBN ???? LCCN ????

Chernoff:1998:FPD

[CHH^{+}98] Anton Chernoff, Mark Herdeg, Ray Hookway, Chris Reeve, Norman Rubin, Tony Tye, S. Bharadwaj Yadavalli, and John Yates. FX!32: A profile-directed binary translator. IEEE Micro, 18(2):56– 64, March/April 1998. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL hf=0?target=
                                        if(eq(query(%27%3CFNO%3E+
                                        cont+m2056%27),0),1,ancestor(
                                        ARTICLE,query(%27%3CFNO%3E+
                                        cont+m2056%27)));   http:/
                                        /dlib.computer.org/dynaweb/
                                        mi/mi1998/@ebt-link;   http:
                                        //dlib.computer.org/mi/books/
                                        mi1998/pdf/m2056.pdf;   http:
                                        //www.computer.org/micro/
                                        mi1998/m2056abs.htm. Pre-

sented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Choquette:1998:GM

[Cho98] Jack Choquette. Genesis microprocessor. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Choquette:2023:NHH

[Cho23] Jack Choquette. NVIDIA Hopper H100 GPU: Scaling performance. IEEE Micro, 43(3): 9–17, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Christie:1995:AKM

[Chr95] D. Christie. AMD-K5 microprocessor. In IEEE [IEE95], pages 41–48. ISBN ???? LCCN ????

Christie:1996:DAK

[Chr96] Dave Christie. Developing the AMD-K5 architecture: Flying without instruments: the independent development on the x86 processor. IEEE Micro, 16(2):16–26, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Conway:2010:CHM

[CKD^{+}10] Pat Conway, Nathan Kalyanasundharam, Gregg Donley, Kevin Lepak, and Bill Hughes. Cache hierarchy and memory subsystem of the AMD Opteron processor. IEEE Micro, 30(2): 16–29, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cismas:2000:AHP

[CM00] Sorin C. Cismas and Kristan J.

Monsen. Architecting highperformance SoC video processors. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Chin:2017:HC

[CM17] Bryan Chin and Subhasish Mitra. Hot Chips 28. IEEE Micro, 37(2):5–6, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020005.html.

Crawford:1990:ICE

[Cra90] John H. Crawford. The i486 CPU: executing instructions in one clock cycle. IEEE Micro, 10(1):27–36, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Chang:2000:MME

[CRTI00] David C. W. Chang, Jayan Ramankutty, Gary Tsztoo, and Sridhar Sharma Isukapalli. MxP — media express processor for VoIP networks. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Cheikh:2021:KDV

[CSM^{+}21] A. Cheikh, S. Sordillo, A. Mastrandrea, F. Menichelli, G. Scotti, and M. Olivieri. Klessydra-T: Designing vector coprocessors for multithreaded edgecomputing cores. IEEE Micro, 41(2):64–71, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cummings:2004:PCC

[Cum04] Uri Cummings. PivotPoint: Clockless crossbar switch for high-performance embedded systems. IEEE Micro, 24 (2):48–59, March/April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2004/02/m2048abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2048.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2048.
pdf.

Chuang:2000:TQC

[CVS^{+}00] Isaac Chuang, Lieven Vandersypen, Matthias Steffan, Gregory Breyta, Costantino Yannoni, and Richard Cleve. Towards quantum computation: a 215hz 5-qubit quantum processor. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Diamant:2023:SIH

[DA23] Ron Diamant and Krste Asanovic.

Special issue on Hot Chips 34. IEEE Micro, 43(3):7–8, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Davies:2002:DMI

[Dav02] David H. Davies. DataPlay’s mobile information distribution and storage technology. IEEE Micro, 22(2):8–15, March/ April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2002/pdf/m2008.
pdf;   http://www.computer.
org/micro/mi2002/m2008abs.
htm.

Douglass:2006:VNG

[DAV06] Steve Douglass, Peter Alfke, and Kees Vissers. Virtex5, the next generation 65nm FPGA. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.S4/HC18.S4T1.
pdf.

Dally:2005:HCP

[DD05] Bill Dally and Keith Diefendorff. Hot Chips 16: Power, parallelism, and memory performance. IEEE Micro, 25(2): 8–9, March/April 2005. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

                                        URL http://csdl.computer.
                                        org/comp/mags/mi/2005/02/
                                        m2008.pdf;   http://csdl.
                                        computer.org/comp/mags/mi/
                                        2005/02/m2008abs.htm.

Diefendorff:1998:ATA

[DDC^{+}98] K. Diefendorff, P. Dubey, R. Chochsprung, et al. AltiVec technology: Accelerating media processing across the spectrum. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Dreslinski:2013:CCS

[DFG^{+}13] Ronald G. Dreslinski, David Fick, Bharan Giridhar, Gyouho Kim, Sangwon Seo, Matthew Fojtik, Sudhir Satpathy, Yoonmyung Lee, Daeyeon Kim, Nurrachman Liu, Michael Wieckowski, Gregory Chen, Dennis Sylvester, David Blaauw, and Trevor Mudge. Centip3De: A 64-core, 3D stacked nearthreshold system. IEEE Micro, 33(2):8–16, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Deaver:1999:WRL

[DGR99] Dean Deaver, Rick Gorton, and Norm Rubin. Wiggins/ Redstone: An on-line program specializer. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

DeHon:1997:MRC

[DHM97] André DeHon, Dan Harman, and Ethan Mirsky. MATRIX: A reconfigurable computing device with configurable instruction distribution. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Ditzel:2000:TCC

[Dit00] David R. Ditzel. Transmeta’s Crusoe: Cool chips for mobile computing. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Djabbari:1996:CVC

[Dja96] Ali Djabbari. Custom VLSI for the compositing DAC of the Touchstone Multimedia Accelerator. In IEEE [IEE96], pages 227–240. ISBN ???? LCCN ????

Darley:1990:TFP

[DKB^{+}90] Merrick Darley, Bill Kronlage, David Bural, Bob Churchill, David Pulling, Paul Wang, Rick Iwamoto, and Larry Yang. The TMS390C602A floatingpoint coprocessor for Sparc systems. IEEE Micro, 10(3): 36–47, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Doweck:2017:IGI

[DKyL^{+}17] Jack Doweck, Wen-Fu Kao, Allen Kuan yu Lu, Julius Mandelblat, Anirudha Rahatekar, Lihu Rappoport, Efraim Rotem, Ahmad Yasin, and Adi Yoaz. Inside 6thgeneration Intel core: New microarchitecture code-named Skylake. IEEE Micro, 37(2): 52–62, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020052-abs.
html.

Dean:2018:NGA

[DPY18] Jeff Dean, David Patterson, and Cliff Young. A new golden age in computer architecture: Empowering the machine-learning revolution. IEEE Micro, 38 (2):21–29, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020021-abs.
html.

Dastidar:2023:AGA

[DRM^{+}23] Jaideep Dastidar, David Riddoch, Jason Moore, Steven Pope, and Jim Wesselkamper. The AMD 400-G adaptive SmartNIC system on chip: a technology preview. IEEE Micro, 43(3):40–49, May/June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Dally:1992:MDP

[DSK^{+}92] William J. Dally, J. A. Stuart Fiske, John S. Keen, Richard A. Lethin, Michael D. Noakes, Peter R. Nuth, Roy E. Davison, and Gregory A. Fyler. The message-driven processor — A multicomputer processing node with efficient mechanisms. IEEE Micro, 12(2):23– 39, March/April 1992. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips III, Stanford University, 1992.

Dally:2001:GEI

[DTB01] William J. Dally, Marc Tremblay, and Allen J. Baum. Guest Editors’ introduction: Hot Chips 12. IEEE Micro, 21(2):13–15, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2001/pdf/m2013.
pdf.

Dubey:1997:ATA

[Dub97] Pradeep Dubey. Afternoon tutorial: Architectural and design implications of mediaprocessing. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Davidson:2018:COS

[DXT^{+}18] Scott Davidson, Shaolin Xie, Christopher Torng, Khalid Al-Hawai, Austin Rovinski, Tutu Ajayi, Luis Vega, Chun Zhao, Ritchie Zhao, Steve Dai, Aporva Amarnath, Bandhav Veluri, Paul Gao, Anuj Rao, Gai Liu, Rajesh K. Gupta, Zhiru Zhang, Ronald Dreslinski, Christopher Batten, and Michael Bedford Taylor. The Celerity open-source 511-core RISC-V tiered accelerator fabric: Fast architectures and design methodologies for fast chips. IEEE Micro, 38(2):30– 41, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020030-abs.
html.

Eddington:2002:IIC

[Edd02] Chris Eddington. InfiniBridge: An InfiniBand channel adapter with integrated switch. IEEE Micro, 22(2):48–56, March/ April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2002/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi2002/m2048abs.
htm.

Eeckhout:2015:HCI

[Eec15] Lieven Eeckhout. Hot chips in an increasingly diverse microprocessor landscape. IEEE Micro, 35(2):2–3, March/April 2015. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://www.

computer.org/csdl/mags/mi/
2015/02/mmi2015020002-abs.
html.

Eeckhout:2016:HCA

[Eec16] Lieven Eeckhout. Hot chips:

The annual feast of riches. IEEE Micro, 36(2):4, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://www.

computer.org/csdl/mags/mi/
2016/02/mmi2016020004-abs.
html.

Eeckhout:2017:HCI

[Eec17] Lieven Eeckhout. Hot chips: Industry and academia cuttingedge microprocessors. IEEE Micro, 37(2):4, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

https://www.computer.org/
csdl/mags/mi/2017/02/mmi2017020004.
html.

Eeckhout:2018:HC

[Eec18] Lieven Eeckhout. Hot Chips 29. IEEE Micro, 38(2):6– 7, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL https://www.computer.
org/csdl/mags/mi/2018/02/
mmi2018020006.html.

Eerola:1997:PRT

[Eer97] Ville Eerola. Pyramid3D: Realtime graphics processor. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Essen:1995:PES

[EG95] A. Essen and S. Goldstein. Performance evaluation of the superscalar speculative execution HaL SPARC64 processor. In IEEE [IEE95], pages 59–74. ISBN ???? LCCN ????

Edenfield:1990:PPM

[EGL^{+}90a] Robin W. Edenfield, Michael G. Gallup, William B. Ledbetter, Jr., Ralph C. McGarity, Eric E. Quintana, and Russel A. Reininger. The 68040 processor: Part 2, memory design and chip verification. IEEE Micro, 10(3): 22–35, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Edenfield:1990:PPD

[EGL^{+}90b] Robin W. Edenfield, Michael G. Gallup, William B. Ledbetter, Jr., Ralph C. McGarity, Eric E. Quintana, and Russell A. Reininger. The 68040 processor: Part I, design and implementation. IEEE Micro, 10(1):66–78, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Essen:1999:BBS

[EM99] Andrew Essen and James Mannos. Broadcom BCM5600 StrataSwitch: A highly integrated Ethernet switch on A chip. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Edmondson:1995:SIE

[ERPR95] John H. Edmondson, Paul Rubinfeld, Ronald Preston, and Vidya Rajagopalan. Superscalar instruction execution in the 21164 Alpha microprocessor. IEEE Micro, 15(2):33– 43, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Einstein:1999:TFI

[ES99] Dan Einstein and Jon Swartz.

TechPress forum: Introduction. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Eberle:2005:ANG

[ESG^{+}05] Hans Eberle, Sheueling Shantz, Vipul Gupta, Nils Gura, Leonard Rarick, and Lawrence Spracklen. Accelerating nextgeneration public-key cryptosystems on general-purpose CPUs. IEEE Micro, 25 (2):52–59, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2005/02/m2052abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2005/02/m2052.
pdf.

Faanes:1998:CVP

[Faa98] Greg Faanes. A CMOS vector processor with a custom streaming cache. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Fang:1999:CTI

[Fan99a] Jesse Fang. Compiler technology on IA-64. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

Fang:1999:DCT

[Fan99b] Jesse Fang. Dynamic compilation technology on IA-64. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Floyd:2011:IAE

[FAWR^{+}11] Michael Floyd, Malcolm Allen-Ware, Karthick Rajamani, Bishop Brock, Charles Lefurgy, Alan J. Drake, Lorena Pesantez, Tilman Gloekler, Jose A. Tierno, Pradip Bose, and Alper Buyuktosunoglu. Introducing the adaptive energy management features of the Power7 chip. IEEE Micro, 31(2):60– 75, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Cordero:1999:SWP

[FCD^{+}99] Frank Ferraiolo, Edgar Cordero, Daniel Dreps, Michael Floyd, Kevin Gower, and Bradley Mc-Credie. A synchronous wavepipeline interface for POWER4. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Flynn:2004:GEI

[FD04] Michael Flynn and Pradeep Dubey. Guest Editors’ introduction: Hot Chips 15— scaling the silicon mountain. IEEE Micro, 24(2):7–9, March/ April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2004/02/m2007.pdf;
http://csdl.computer.org/
dl/mags/mi/2004/02/m2007.
htm.

Foley:2017:UPP

[FD17] Denis Foley and John Danskin. Ultra-performance Pascal GPU and NVLink interconnect. IEEE Micro, 37(2): 7–17, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020007-abs.
html.

Frank:1999:BBC

[FH99] Ed Frank and Jack Holloway. BCM4100/BCM4210: A chipset for 16 Mbps phoneline networking. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Frank:2000:CHP

[FH00] Edward H. Frank and Jack Holloway. Connecting the home with a phone line network chip set. IEEE Micro, 20(2):27–38, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m2027.
pdf;   http://www.computer.
org/micro/mi2000/m2027abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Feehrer:2013:OST

[FJL^{+}13] John Feehrer, Sumti Jairath, Paul Loewenstein, Ram Sivaramakrishnan, David Smentek, Sebastian Turullols, and Ali Vahidsafa. The Oracle Sparc T5 16-core processor scales to eight sockets. IEEE Micro, 33(2): 48–57, March/April 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Fleischmann:2000:CPM

[Fle00] Marc Fleischmann. Crusoe power management: Cutting x86 operating power through LongRun. In IEEE [IEE00], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/index12.html. Best pre-

sentation award.

Fandrianto:1995:SCV

[FM95] J. Fandrianto and B. Martin. A single chip video CD with hifi audio for consumer applications. In IEEE [IEE95], pages 135–142. ISBN ???? LCCN ????

Farley:2018:APN

[FME18] Brendan Farley, John McGrath, and Christophe Erdmann. An all-programmable 16-nm RF-SoC for digital-RF communications. IEEE Micro, 38(2): 61–71, March/April 2018. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020061-abs.
html.

Fossum:1998:DCS

[Fos98] Eric R. Fossum. Digital camera system on a chip. IEEE Micro, 18(3):8–15, May/June 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL hf=

0?target=if(eq(query(%27%
3CFNO%3E+cont+m3008%27),0)
,1,ancestor(ARTICLE,query(
%27%3CFNO%3E+cont+m3008%27)
));   http://dlib.computer.
org/dynaweb/mi/mi1998/@ebt-
link;   http://dlib.computer.
org/mi/books/mi1998/pdf/m3008.
pdf;   http://www.computer.
org/micro/mi1998/m3008abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Fisch:2006:ZRU

[FSP06] David Fisch, Anant Singh, and Greg Popov. Z-RAM ultra-dense memory for 90nm and below. Web document., 2006. URL http://

www.hotchips.org/archives/

                                        hc18/2_Mon/HC18.S3/HC18.S3T1.
                                        pdf.

Fan:2012:GEM

[FZW^{+}12] Dongrui Fan, Hao Zhang, Da Wang, Xiaochun Ye, Fenglong Song, Guojie Li, and Ninghui Sun. Godson-T: An efficient many-core processor exploring thread-level parallelism. IEEE Micro, 32(2):38– 47, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Grossman:2021:XSX

[GA21] M. Grossman and J. Andrews. The Xbox Series X system architecture. IEEE Micro, 41(2): 22–28, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gandhi:1998:SMR

[Gan98] Prashant P. Gandhi. SA-1500: A 300 MHz RISC CPU with attached media processor. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Garibay:1995:BBB

[Gar95] T. Garibay. Building a better beast: Native vs. RISC-like vs. VLIW methods of implementing x86 microprocessors. In IEEE [IEE95], pages 49–58. ISBN ???? LCCN ????

Garside:1996:A

[Gar96] Jim Garside. AMULET2e. In IEEE [IEE96], pages 257–274. ISBN ???? LCCN ????

Garside:2000:AAS

[Gar00] Jim Garside. AMULET3i: An asynchronous system-on-chip. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http:/

/www.hotchips.org/index12.
html.

Grisenthwaite:2023:AME

[GBW^{+}23] Richard Grisenthwaite, Graeme Barnes, Robert N. M. Watson, Simon W. Moore, Peter Sewell, and Jonathan Woodruff. The Arm Morello Evaluation Platform validating CHERI-based security in a high-performance system. IEEE Micro, 43(3): 50–57, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gilbert:2008:GUW

[GDES08] Jeffrey M. Gilbert, Chinh H.

Doan, Sohrab Emami, and C. Bernard Shung. A 4-Gbps uncompressed wireless HD A/V transceiver chipset. IEEE Micro, 28(2):56–64, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Goulding-Hotta:2011:GMA

[GHSV^{+}11] Nathan Goulding-Hotta, Jack Sampson, Ganesh Venkatesh, Saturnino Garcia, Joe Auricchio, Po-Chao Huang, Manish Arora, Siddhartha Nath, Vikram Bhatt, Jonathan Babb, Steven Swanson, and Michael Bedford Taylor. The GreenDroid Mobile Application Processor: An architecture for silicon’s dark future. IEEE Micro, 31(2): 86–95, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Guo:2017:SHC

[GHY^{+}17] Kaiyuan Guo, Song Han, Song Yao, Yu Wang, Yuan Xie, and Huazhong Yang. Softwarehardware codesign for efficient neural network acceleration. IEEE Micro, 37(2):18– 25, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/mi/
                                        2017/02/mmi2017020018-abs.
                                        html.

Galles:2021:PDS

[GM21] M. Galles and F. Matus. Pensando distributed services architecture. IEEE Micro, 41(2): 43–49, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Golston:2000:TAE

[Gol00] Jeremiah Golston. TMS320 C64x architecture extensions boost performance for broadband communications and imaging. In IEEE [IEE00], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/index12.html.

Gonzalez:1999:CEP

[Gon99] Ricardo E. Gonzalez. Configurable and extensible processors change system design. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Gonzalez:2000:XCE

[Gon00] Ricardo E. Gonzalez. Xtensa — A configurable and extensible processor. IEEE Micro, 20(2):60–70, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m2060.
pdf;   http://www.computer.
org/micro/mi2000/m2060abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Greenstein:2011:DBS

[Gre11] Shane Greenstein. The direction of broadband spillovers. IEEE Micro, 31(2):104, 103, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Gillmore:1999:TFP

[GTB99] Dan Gillmore, Dean Takahasi, and Matt Beers. TechPress forum: Panel discussion: For the media. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Hayes:1997:RCP

[Hay97] Ken Hayes. Reality co-processor — the power in Nintendo64. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Holzle:1997:JSS

[HBG^{+}97] Urs Hölzle, Lars Bak, Steffen Grarup, Robert Griesemer, and Srdjan Mitrovic. Java on steroids: Sun’s highperformance Java implementation. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Heddes:2000:PNP

[Hed00] Marco Heddes. Power network processor architecture. In IEEE [IEE00], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/index12.html.

Hester:2007:KIM

[Hes07] Phil Hester. Keynote II: Multicore and beyond: Evolving the x86 architecture. Web document., July/August 21, 2007.

                                        URL http://www.hotchips.
                                        org/hc19/docs/keynote2.pdf.

Hsu:1998:DSC

[hH98] Feng hsiung Hsu. Designing a single chip chess grandmaster while knowing nothing about chess [well, I did know close to nothing about chess…]. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Hammond:1999:SHC

[HHS^{+}99] Lance Hammond, Ben Hubbert, Michael Siu, Manohar Prabhu, Mark Willey, Michael Chen, Maciek Kozyrczak, and Kunle Olukotun. The Stanford Hydra CMP. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

Hammond:2000:SHC

[HHS^{+}00] Lance Hammond, Benedict A.

Hubbert, Michael Siu, Manohar K. Prabhu, Michael Chen, and Kunle Olukotun. The Stanford Hydra CMP. IEEE Micro, 20(2):71–84, March/ April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m2071.
pdf;   http://www.computer.
org/micro/mi2000/m2071abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Hammarlund:2014:HFG

[HMB^{+}14] Per Hammarlund, Alberto J.

Martinez, Atiq A. Bajwa, David L. Hill, Erik Hallnor, Hong Jiang, Martin Dixon, Michael Derr, Mikal Hunsaker, Rajesh Kumar, Randy B. Osborne, Ravi Rajwar, Ronak Singhal, Reynold D’Sa, Robert Chappell, Shiv Kaushik, Srinivas Chennupaty, Stephan Jourdan, Steve Gunther, Tom Piazza, and Ted Burton. Haswell: The fourth-generation Intel core processor. IEEE Micro, 34(2): 6–20, March/April 2014. CO-DEN IEMIDZ. ISSN 0272-1732.

Hohensee:1996:WCE

[HMR96] Paul Hohensee, Mat Myszewski, and David Reese. The Wabi CPU emulator technology. In IEEE [IEE96], pages 47–66. ISBN ???? LCCN ????

Haring:2012:IBG

[HOF^{+}12] Ruud A. Haring, Martin Ohmacht, Thomas W. Fox, Michael K. Gschwind, David L. Satterfield, Krishnan Sugavanam, Paul W. Coteus, Philip Heidelberger, Matthias A. Blumrich, Robert W. Wisniewski, Alan Gara, George Liang-Tai Chiu, Peter A. Boyle, Norman H. Chist, and Changhoan Kim. The IBM Blue Gene/Q compute chip. IEEE Micro, 32(2):48–60, March/ April 2012. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hootman:1990:HC

[Hoo90] J. Hootman. Hot Chips. 2. IEEE Micro, 10(3):3–4, May/ June 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hundt:1997:HOA

[Hun97]  Reed Hundt.  HDTV and

other advances in communications and broadcasting. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Hill:1991:GEI

[HW91] Mark D. Hill and David A. Wood. Guest Editors’ introduction: Hot chips II symposium. IEEE Micro, 11(3):8–9, May/ June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Hu:2009:GSM

[HWG^{+}09] Weiwu Hu, Jian Wang, Xiang Gao, Yunji Chen, Qi Liu, and Guojie Li. Godson-3: A scalable multicore RISC processor with x86 emulation. IEEE Micro, 29(2):17–29, March/April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Holmann:1996:VPM

[HYYS96] Edgar Holmann, Toyohiko Yoshida, Akira Yamada, and Yukihiko Shimazu. A VLIW processor for multimedia applications. In IEEE [IEE96], pages 193–202. ISBN ???? LCCN ????

Iyer:2000:CCN

[IDTS00] Sundar Iyer, Ajay Desai, Ajay Tambe, and Ajit Shelat. ClassiPI: A classifier for next generation content and policy based switches. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

IEEE:1993:HCV

[IEE93]  IEEE, editor.  Hot chips

V, a symposium on highperformance chips: Stanford University, CA, August 8–10, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN ???? LCCN ????

IEEE:1994:HCV

[IEE94] IEEE, editor. Hot Chips VI, a symposium on high performance chips: Stanford University, CA, August 14–16, 1994. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN ???? LCCN ????

IEEE:1995:HCV

[IEE95] IEEE, editor. Hot chips VII: symposium record: Stanford University, Stanford, California, August 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? LCCN ????

IEEE:1996:HCV

[IEE96] IEEE, editor. Hot chips VIII: symposium record: Stanford University, Stanford, California, August 18–20, 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? LCCN ????

IEEE:1997:HCI

[IEE97] IEEE, editor. Hot Chips IX: Stanford University, Stanford, California, August 24–26, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN ???? LCCN ????

IEEE:1998:HCC

[IEE98] IEEE, editor. Hot chips 10: conference record: August 16– 18, 1998, Memorial Auditorium, Stanford University, Palo Alto, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ????

IEEE:1999:HCS

[IEE99] IEEE, editor. Hot Chips 11:

Stanford University, Stanford, California, August 15–17, 1999. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

IEEE:2000:HCS

[IEE00] IEEE, editor. Hot Chips 12:

Stanford University, Stanford, California, August 13–15, 2000. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

IEEE:2001:HCS

[IEE01] IEEE, editor. Hot Chips 13:

Stanford University, Stanford, California, August 19–21, 2001, volume 22(2) of IEEE Micro. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN ???? LCCN

????  URL http://www.
hotchips.org/hc13/.

IEEE:2002:HCS

[IEE02] IEEE, editor. Hot Chips 14:

Stanford University, Stanford, California, August 18–20, 2002. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/hc14/.

IEEE:2003:HCS

[IEE03] IEEE, editor. Hot Chips 15: Stanford University, Stanford, California, August 17–19, 2003. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/hc15/.

IEEE:2004:HCS

[IEE04] IEEE, editor. Hot Chips 16: Stanford University, Stanford, California, August 22–24, 2004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/hc16/.

IEEE:2005:HCS

[IEE05] IEEE, editor. Hot Chips 17: Stanford University, Stanford, California, August 14–16, 2005. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/hc17/.

IEEE:2006:HCS

[IEE06] IEEE, editor. Hot Chips 18:

Stanford University, Stanford, California, August 20–22, 2006. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2006. ISBN ???? LCCN

????  URL http://www.
hotchips.org/hc18/.

IEEE:2007:HCS

[IEE07] IEEE, editor. Hot Chips 19:

Stanford University, Stanford, California, August 19–21, 2007. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2007. ISBN ???? LCCN

????  URL http://www.
hotchips.org/hc19/.

IEEE:2008:HCS

[IEE08] IEEE, editor. Hot Chips 20:

Stanford University, Stanford, California, August 24–26, 2008. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2008. ISBN ???? LCCN

????  URL http://www.
hotchips.org/hc20/.

IEEE:2009:HCS

[IEE09] IEEE, editor. Hot Chips 21:

Stanford University, Stanford, California, August 23–25, 2009. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2009. ISBN ???? LCCN

????  URL http://www.
hotchips.org/hc21/.

IEEE:2010:HCS

[IEE10] IEEE, editor. Hot Chips 22: Stanford University, Stanford, California, August 15–17, 2010. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2010. ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/hc22/.

IEEE:2011:HCS

[IEE11] IEEE, editor. Hot Chips 23: Stanford University, Stanford, California, August 17–19, 2011. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN ???? LCCN

                                        ???? URL http://hotchips.
                                        org/archives/hot-chips-23.

IEEE:2012:HCS

[IEE12] IEEE, editor. Hot Chips 24: Flint Center, Cupertino, California, August 27–29, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN ???? LCCN

                                        ???? URL http://hotchips.
                                        org/archives/hot-chips-24.

IEEE:2013:HCS

[IEE13] IEEE, editor. Hot Chips 25: Stanford Memorial Auditorium, Stanford, California, August 25–27, 2013. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2013. ISBN ???? LCCN ???? URL http://

hotchips.org/archives/hot-
chips-24.

Jouppi:1996:GEI

[JA96] Norman P. Jouppi and Hasan S.

Alkhatib. Guest Editors’ introduction: Hot chips and the microprocessor. IEEE Micro, 16 (2):6–7, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

James:1990:MBE

[Jam90] David V. James. Multiplexed buses — the endian wars continue. IEEE Micro, 10(3): 9–21, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jayavant:1998:MHP

[Jay98] Rajeev Jayavant. MXi: A high-performance x86 processor with integrated 3D graphics. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Johnson:1990:HCS

[Joh90]  Stephen C. Johnson.  Hot

chips and soggy software: RISC success springs partially from good system design. take note and eliminate the software bottleneck from your new design. IEEE Micro, 10(1):23–26, January/February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Johnson:1998:TMM

[Joh98] David Johnson. Techniques for mitigating memory latency effects in the PA-8500 processor.

In IEEE [IEE98], page ?? ISBN ???? LCCN ????

John:2019:EHC

[Joh19] L. John. Emerging hot chips and systems. IEEE Micro, 39 (2):4–5, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2020:DMC

[Joh20] L. K. John. Did ML chips heat up the chip design arena? IEEE Micro, 40(2):4–5, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2022:HCM

[Joh22] Lizy Kurian John. Hot Chips 33 and more! IEEE Micro, 42 (3):4–5, May/June 2022. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

John:2023:HCM

[Joh23] Lizy Kurian John. Hot Chips 34 and more! IEEE Micro, 43 (3):4–6, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Jouppi:1992:HCI

[Jou92] Norman P. Jouppi. Hot Chips-III — introduction. IEEE Micro, 12(2):8–9, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Joy:1996:MAL

[Joy96] Bill Joy. Microprocessor architecture: Looking forward. In IEEE [IEE96], page ?? ISBN ???? LCCN ????

Jouppi:1999:PIA

[JSR^{+}99] Norm Jouppi, Ken Shoemaker, Kathy Richardson, David Armitage, Jim Barton, Natasha Flaherty, and Elizabeth Houck. Panel: Information appliances in the home. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Jouppi:1998:MPC

[JW98] Norman P. Jouppi and John Wawrzynek. Message from the program co-chairs. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Kagan:1996:PMF

[Kag96] Michael Kagan. The P55C microarchitecture — the first implementation of MMX technology. In IEEE [IEE96], pages 157–162. ISBN ???? LCCN ????

Kalapathy:1996:HSI

[Kal96] Paul Kalapathy. Hardware/ software interaction on the Mpact media processor. In IEEE [IEE96], pages 179–192. ISBN ???? LCCN ????

Kongetira:2005:NWM

[KAO05] Poonacha Kongetira, Kathirgamar Aingaran, and Kunle Olukotun. Niagara: A 32way multithreaded Sparc processor. IEEE Micro, 25 (2):21–29, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2005/02/m2021abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2005/02/m2021.
pdf.

Kozyrakis:2020:HCR

[KB20] C. Kozyrakis and I. Bratt. The hot chips renaissance. IEEE Micro, 40(2):6–7, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krishnan:2016:EEG

[KBN16] Guhan Krishnan, Dan Bouvier, and Samuel Naffziger. Energyefficient graphics and multimedia in 28-nm Carrizo accelerated processing unit. IEEE Micro, 36(2):22–33, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http://www.

computer.org/csdl/mags/mi/
2016/02/mmi2016020022-abs.
html.

Khailany:2001:IMP

[KDK^{+}01] Brucek Khailany, William J. Dally, Ujval J. Kapasi, Peter Mattson, Jinyung Namkoong, John D. Owens, Brian Towles, Andrew Chang, and Scott Rixner. Imagine: Media processing with streams. IEEE Micro, 21(2):35–46, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/

mi/books/mi2001/pdf/m2035.
pdf;   http://www.computer.
org/micro/mi2001/m2035abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Khailany:2000:ISI

[KDR^{+}00] Brucek Khailany, William J.

Dally, Scott Rixner, Ujval J. Kapasi, Peter Mattson, Jinyung Namkoong, John D. Owens, and Brian Towles. IMAGINE: Signal and image processing using streams. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://cva.
stanford.edu/imagine.

Keckler:1997:MC

[Kec97] Steve Keckler. The MIT MAP chip. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Kessler:1998:AMO

[Kes98] Richard E. Kessler. The Alpha 21264 microprocessor: Out-of-order execution at 600 MHz. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Knies:1999:TIA

[KFL99] Allan Knies, Jesse Fang, and Wei Li. Tutorial: IA64 architecture and compilers. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Kozyrakis:2000:VIM

[KGM^{+}00] Christoforos Kozyrakis, Joseph Gebis, David Martin, Samuel Williams, Ioannis Mavroidis, Steven Pope, Darren Jones, and David Patterson. Vector IRAM: A media-enhanced vector processor with embedded DRAM. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Khurana:1996:BWG

[Khu96] A. Khurana. Bringing workstation graphics performance to a desktop near you: ViRGE VX. In IEEE [IEE96], pages 289– 298. ISBN ???? LCCN ????

Kidd:2014:PCO

[Kid14]  David Kidd.  Process and

circuit optimization for power reduction using DDC transistors. IEEE Micro, 34(2):54– 62, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Kunimatsu:1999:GVU

[KIS^{+}99] A. Kunimatsu, N. Ide, T. Sato, Y. Endo, H. Murakami, T. Kamei, M. Hirano, M. Oka, A. Ophba, T. Yutaka, T. Okada, and M. Suzuoki. 5.5 GFLOPS vector units for “emotion synthesis”. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Kunimatsu:2000:VUA

[KIS^{+}00] Atsushi Kunimatsu, Nobuhiro Ide, Toshinori Sato, Yukio Endo, Hiroaki Murakami, Takayuki Kamei, Masashi Hirano, Fujio Ishihara, Haruyuki Tago, Masaaki Oka, Akio Ohba, Teiji Yutaka, Toyoshi Okada, and Masakazu Suzuoki. Vector unit architecture for emotion synthesis. IEEE Micro, 20(2):40–47, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m2040.
pdf;   http://www.computer.
org/micro/mi2000/m2040abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Keltcher:1999:AAN

[KKK^{+}99] Chetana Keltcher, Jim Kelly, Ramani Krishnan, John Peck, Steve Polzin, Sr.idhar Subramanian, and Fred Weber. AMD Athlon Northbridge with 4x AGP and next generation memory subsystem. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Klecha:2006:HEQ

[KKO06] Marcin Klecha, Ralf Karge, and Richard O’Connor. Home entertainment-quality multimedia experience whilst on the move — Philips Nexperia Mobile Multimedia Co-Processor PNX4103. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.S1/HC18.S1T3.
pdf.

Kurihara:1999:RPS

[KKSS99] Toshihiko Kurihara, Eiki Kamada, Kentaro Shimada, and Teruhisa Shimizu. A RISC processor for SR8000: Accelerating large scale scientific computing with SMP. In IEEE [IEE99], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/hotc11_index.
                                        html.

Koufaty:2003:HTN

[KM03] David Koufaty and Deborah T. Marr. Hyperthreading technology in the netburst microarchitecture. IEEE Micro, 23(2):56–65, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2003/pdf/m2056.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m2056abs.
                                        htm.

Keltcher:2003:AOP

[KMAC03] Chetana N. Keltcher, Kevin J. McGrath, Ardsher Ahmed, and Pat Conway. The AMD Opteron processor for multiprocessor servers. IEEE Micro, 23(2):66–76, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2003/pdf/m2066.
                                        pdf;   http://www.computer.
                                        org/micro/mi2003/m2066abs.
                                        htm.

Kapil:2004:CMP

[KML04] Sanjiv Kapil, Harlan McGhan, and Jesse Lawrendra. A chip multithreaded processor for network-facing workloads. IEEE Micro, 24 (2):20–30, March/April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2004/02/m2020abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2020.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2020.
pdf.

Knies:1999:IAB

[Kni99a] Allan Knies. IA-64 architecture basics/introduction. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Knies:1999:OTU

[Kni99b] Allan Knies. Optimization techniques/using IA-64 features. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Kota:2005:HLS

[KO05] Rajesh Kota and Rich Oehler.

Horus: Large-scale symmetric multiprocessing for Opteron systems. IEEE Micro, 25 (2):30–40, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/mi/2005/02/m2030abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2005/02/m2030.
                                        pdf.

Kilgariff:1996:TFA

[KR96] Emmett Kilgariff and Martin Randall. Touchstone — A fresh approach to multimedia for the PC. In IEEE [IEE96], pages 203–216. ISBN ???? LCCN ????

Kubiatowicz:2019:HC

[KR19] J. Kubiatowicz and S. Rusu. Hot Chips 30. IEEE Micro, 39 (2):6–8, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Krech:1998:BLS

[Kre98] Alan Krech. Blitzen: Lightning speed 3D geometry accelerator. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Kruckemyer:2000:SCH

[Kru00] David Kruckemyer. The SB-1 core: A high-performance, low-power MIPS64 implementation. In IEEE [IEE00], page ?? ISBN ???? LCCN

                                        ????  URL http://www.
                                        hotchips.org/index12.html.

Kitahara:1990:GBM

[KS90] Takeshi Kitahara and Taizo Satoh. The Gmicro/300 32-Bit microprocessor. IEEE Micro, 10 (3):68–75, May/June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kubiatowicz:2007:GEI

[KS07] John Kubiatowicz and Howard Sachs. Guest Editors’ introduction: Hot Chips 18. IEEE Micro, 27(2):7–9, March/ April 2007. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2007/02/m2007.pdf.

Kondo:1996:TCM

[KSI^{+}96] Toshio Kondo, Kazuhito Suguri, Mitsuo Ikeda, Tetsuya Abe, Hiroaki Matsuda, Tsuneo Okubo, Kenji Ogura, Yutaka Tashiro, Naoki Ono, Toshihiro Minami, Ritsu Kusaba, Takeshi Ikenaga, Nobutaro Shibata, Ryota Kasai, Koji Otsu, Fumiaki Nakagawa, and Yasuhiko Sato. Two-chip MPEG-2 video encoder: Switching to simple profile at main level for a costeffective encoder. IEEE Micro, 16(2):51–58, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Kondo:1995:TCR

[KSIA95] T. Kondo, K. Suguri, M. Ikeda, and T. Abe. A two-chip real-time MPEG2 video encoder with wide range motion estimation. In IEEE [IEE95], pages 95–102. ISBN ???? LCCN ????

Kalla:2010:PIN

[KSSF10] Ron Kalla, Balaram Sinharoy, William J. Starke, and Michael Floyd. Power7: IBM’s next-generation server processor. IEEE Micro, 30(2):7– 15, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kalla:2004:IPC

[KST04] Ron Kalla, Balaram Sinharoy, and Joel M. Tendler. IBM Power5 chip: A dual-core multithreaded processor. IEEE Micro, 24(2):40–47, March/ April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/mi/2004/02/m2040abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/02/m2040.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2004/02/m2040.
                                        pdf.

Knight:1999:SIA

[KTP^{+}99] J. Knight, R. Tso, L. Peng, A. Pande, and G. Turetzky. SiRFstar II architecture: A powerful system platform for consumer GPS applications. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Kumar:1996:HPR

[Kum96] Ashok Kumar. The HP PA-8000 RISC CPU: A high performance out-of-order processor. In IEEE [IEE96], pages 9–20. ISBN ???? LCCN ????

KurianJohn:2021:CGM

[Kur21] L. Kurian John. CPUs, GPUs, and more from Hot Chips 32. IEEE Micro, 41(2):4–5, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kutaragi:1999:KAN

[Kut99] Ken Kutaragi. Keynote address: New millenium for computer entertainment. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Kozyrakis:2009:HCT

[KvdW09] Christos Kozyrakis and Jan-Willem van de Waerdt. Hot Chips turns 20. IEEE Micro, 29 (2):4–5, March/April 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Kubiatowicz:2002:GEI

[KW02] John Kubiatowicz and Andrew Wolfe. Guest Editors’ introduction: Hot Chips 13. IEEE Micro, 22(2):6–7, March/ April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2002/pdf/m2006.
pdf;   http://www.computer.
org/micro/mi2002/m2006abs.
htm.

Kozyrakis:2013:SRH

[KZ13] Christos Kozyrakis and Rumi Zahir. Selected research from Hot Chips 24. IEEE Micro, 33 (2):6–7, March/April 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Larri:1996:ADB

[Lar96] Guy Larri. ARM810: Dancing to the beat of a different drum. In IEEE [IEE96], pages 109–118. ISBN ???? LCCN ????

Lin:2006:MPT

[LATSK06] Yuan Lin, Ali-Reza Adl-Tabatabai, Bratin Saha, and Christos Kozyrakis. Multicore programming: From threads to transactional memory. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/1_Sun/HC18.T1P1.pdf;
 http://www.hotchips.org/
archives/hc18/1_Sun/HC18.
T1P2.pdf;   http://www.hotchips.
org/archives/hc18/1_Sun/HC18.
T1P3.pdf;   http://www.hotchips.
org/archives/hc18/1_Sun/HC18.
T1P4.pdf.

Lam:2000:GEI

[LB00] Monica Lam and Forest Baskett. Guest Editors’ introduction: Cutting-edge designs. IEEE Micro, 20(2):14– 15, March/April 2000. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL http://dlib.computer.
org/mi/books/mi2000/pdf/m2014.

pdf. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Lindtjorn:2011:BTM

[LCP^{+}11] Olav Lindtjorn, Robert Clapp, Oliver Pell, Haohuan Fu, Michael Flynn, and Oskar Mencer. Beyond traditional microprocessors for geoscience high-performance computing applications. IEEE Micro, 31 (2):41–49, March/April 2011. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lewis:1998:VIA

[LD98] Michael C. Lewis and Joseph C.

Del Rio. VelaTX: Innovative 3D architecture coupled with embedded DRAM architecture. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Lee:1995:AME

[Lee95] Ruby B. Lee. Accelerating multimedia with enhanced microprocessors. IEEE Micro, 15(2): 22–32, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Lee:1997:EMM

[Lee97] Ruby Lee. Effectiveness of the MAX-2 multimedia extensions for PA-RISC 2.0 processors. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Lie:2023:CAD

[Lie23] Sean Lie. Cerebras architecture deep dive: First look inside the hardware/software co-design for deep learning. IEEE Micro, 43 (3):18–30, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lynch:1998:UIM

[LL98] Bill Lynch and Gary Lauterbach. UltraSPARC III: A 600 MHz 64-bit superscalar processor for 1000-way scalable systems. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Lindholm:2008:NTU

[LNOM08] Erik Lindholm, John Nickolls, Stuart Oberman, and John Montrym. NVIDIA Tesla: A unified graphics and computing architecture. IEEE Micro, 28(2):39–55, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Litch:1998:SPC

[LS98] Tim Litch and Jeff Slaton. StrongARMing portable communications. IEEE Micro, 18(2):48–55, March/April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL hf=

0?target=if(eq(query(%27%
3CFNO%3E+cont+m2048%27),0)
,1,ancestor(ARTICLE,query(
%27%3CFNO%3E+cont+m2048%27)
));   http://dlib.computer.
org/dynaweb/mi/mi1998/@ebt-
link;   http://dlib.computer.
org/mi/books/mi1998/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi1998/m2048abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Lucas:1999:HSL

[Luc99] L. Lucas. High speed low cost TM1300 Trimedia enhanced PCI VLIW mediaprocessor. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Litz:2024:SIH

[LV24] Heiner Litz and Natalia Vassilieva. Special issue on Hot Chips 2023. IEEE Micro, 44 (3):6–7, May/June 2024. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Lee:2016:AAB

[LWC^{+}16] Yunsup Lee, Andrew Waterman, Henry Cook, Brian Zimmer, Ben Keller, Alberto Puggelli, Jaehwa Kwak, Ruzica Jevtic, Stevo Bailey, Milovan Blagojevic, Pi-Feng Chiu, Rimas Avizienis, Brian Richards, Jonathan Bachrach, David Patterson, Elad Alon, Bora Nikolic, and Krste Asanovic. An agile approach to building RISC-V microprocessors. IEEE Micro, 36(2):8–20, March/April 2016. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2016/02/mmi2016020008-abs.
html.

Maher:1996:MIS

[Mah96] Robert Maher. Multimedia instruction set extensions for a sixth-generation x86 processor. In IEEE [IEE96], pages 163– 170. ISBN ???? LCCN ????

Malinowski:1997:TCA

[Mal97] Richard Malinowski. Technical challenges associated with the development of the Intel 440LX AGPset. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Mansur:2009:NNF

[Man09] Dan Mansur. A new 40-nm FPGA and ASIC common platform. IEEE Micro, 29(2):46– 53, March/April 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Mateosian:1997:MNV

[Mat97]  R. M. Mateosian.  Micro

news: Visiting Hot Chips IX. IEEE Micro, 17(5):5, September/October 1997. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL hf=0?target=
                                        if(eq(query(%27%3CFNO%3E+
                                        cont+m5005%27),0),1,ancestor(
                                        ARTICLE,query(%27%3CFNO%3E+
                                        cont+m5005%27)));   http:/
                                        /dlib.computer.org/dynaweb/
                                        mi/mi1997/@ebt-link;   http:
                                        //dlib.computer.org/mi/books/
                                        mi1997/pdf/m5005.pdf.

Mateosian:2011:T

[Mat11] Richard Mateosian. Technology. IEEE Micro, 31(2):100– 102, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McNairy:2005:MDC

[MB05] Cameron McNairy and Rohit Bhatia. Montecito: A dual-core, dual-thread Itanium processor. IEEE Micro, 25 (2):10–20, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2005/02/m2010abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2005/02/m2010.
pdf.

McCredie:1999:PTD

[MBB^{+}99] B. McCredie, J. Badar, R. Bailey, P. Chou, C. Carter, D. Dreps, J. Eckhardt, D. Ervin, M. Floyd, A. Haridass, D. Heidel, M. Immediato, J. Keaty, B. Krauter, J. LeBlanc, L. Leitner, D. Malone, D. Mikan, Jr., M. Nealon, J. Petrovick, D. Plass, K. Reick, P. Restle, R. Robertazzi, T. Skergan, K. Stawiasz, H. Stigdon, J. Vargus, and J. D. Warnock. A 1GHz POWER4 testchip design. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

McConnell:1999:MPS

[McC99] Ray McConnell. Massively parallel SIMD computing on the FUZION chip. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

McMinn:1995:FSF

[McM95] B. McMinn. The first superscalar 29K family member. In IEEE [IEE95], pages 1–10. ISBN ???? LCCN ????

Mutz:2006:HME

[MD06] Stephane Mutz and Philippe Durieux. Heterogeneous multiprocessing for efficient multistandard high definition video decoding. Web document.,

                                        2006.  URL http://www.
                                        hotchips.org/archives/hc18/
                                        2_Mon/HC18.S1/HC18.S1T2.pdf.

Medina:2020:HLP

[MD20] E. Medina and E. Dagan. Habana Labs purpose-built AI inference and training processor architectures: Scaling AI training systems using standard Ethernet with Gaudi processor. IEEE Micro, 40(2):17– 24, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Meindl:1997:GIS

[Mei97] James Meindl. Gigascale integration: Is the sky the limit? In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Meyerson:2006:CIN

[Mey06] Bernard Meyerson. Collaborative innovation and a new lever in information technology development. Web document., July/August 2006.

URL http://www.hotchips.
org/hc18/docs/keynote2_hc18.
pdf.

McKeown:2017:PMP

[MFN^{+}17] Michael McKeown, Yaosheng Fu, Tri Nguyen, Yanqi Zhou, Jonathan Balkind, Alexey Lavrov, Mohammad Shahrad, Samuel Payne, and David Wentzlaff. Piton: A manycore processor for multitenant clouds. IEEE Micro, 37(2):70– 80, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020070-abs.
html.

Minami:1998:SCM

[MKN^{+}98] Toshihiro Minami, T. Kondo, K. Nitta, K. Suguri, M. Ikeda, T. Yoshitome, H. Watanabe, H. Iwasaki, K. Ochiai, J. Naganuma, M. Endo, E. Yamagishi, T. Takahashi, K. Tadaishi, Y. Tashiro, N. Kobayashi, T. Okubo, T. Ogura, and R. Kasai. A single-chip MPEG2 MP@ML video encoder LSI with multi-chip configuration for a single-board MP@HL encoder. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Mattioli:2021:HPW

[ML21] M. Mattioli and A. Lahtiranta.

Hidden potential within video game consoles. IEEE Micro, 41(2):72–77, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Montrym:1996:IGP

[MM96] John Montrym and Brian Mc-Clendon. InfiniteReality Graphics — power through complexity. In IEEE [IEE96], pages 299– 308. ISBN ???? LCCN ????

Montrym:2005:G

[MM05] John Montrym and Henry Moreton. The GeForce 6800. IEEE Micro, 25 (2):41–51, March/April 2005. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

                                        //csdl.computer.org/comp/
                                        mags/mi/2005/02/m2041abs.
                                        htm;   http://csdl.computer.
                                        org/dl/mags/mi/2005/02/m2041.
                                        pdf.

McCormack:1998:NBF

[MMG^{+}98] Joel McCormack, Bob Mc-Namara, Chris Gianos, Larry Seiler, Norm Jouppi, and Ken Correll. Neon: A big, fast, 3D workstation graphics accelerator. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Modi:1997:PIC

[Mod97] Nimish Modi. The PentiumAE II CPU: A high performance dynamic execution processor with MMX technology. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Mattson:2020:MIS

[MRC^{+}20] P. Mattson, V. J. Reddi, C. Cheng, C. Coleman, G. Diamos, D. Kanter, P. Micikevicius, D. Patterson, G. Schmuelling, H. Tang, G. Wei, and C. Wu. MLPerf: An industry standard benchmark suite for machine learning performance. IEEE Micro, 40(2):8–16, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

McNairy:2003:IPM

[MS03] Cameron McNairy and Don Soltis. Itanium 2 processor microarchitecture. IEEE Micro, 23(2):44–55, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2003/pdf/m2044.
pdf;   http://www.computer.
org/micro/mi2003/m2044abs.
htm.

Mirapuri:1992:MRP

[MWV92] Sunil Mirapuri, Michael Woodacre, and Nader Vasseghi. The MIPS R4000 processor. IEEE Micro, 12(2):10–22, March/April 1992. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips III, Stanford University, 1992.

Maruyama:2010:SVN

[MYK^{+}10] Takumi Maruyama, Toshio Yoshida, Ryuji Kan, Iwao Yamazaki, Shuji Yamamura, Noriyuki Takahashi, Mikio Hondou, and Hiroshi Okano. Sparc64 VIIIfx: A newgeneration octocore processor for petascale computing. IEEE Micro, 30(2):30–40, March/ April 2010. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Naas:1995:MPF

[Naa95] B. Naas. Memory performance features of the 64-bit PA-8000. In IEEE [IEE95], pages 87–94. ISBN ???? LCCN ????

Nowatzyk:1995:SMC

[NABR95] A. Nowatzyk, G. Aybay, M. Browne, and B. Radke. Scylla: A memory controller with integrated protocol engines for distributed shared memory support. In IEEE [IEE95], pages 179–186. ISBN ???? LCCN ????

Normoyle:1998:UIE

[NCT^{+}98] Kevin B. Normoyle, Michael A. Csoppenszky, Allan Tzeng, Timothy P. Johnson, Christopher D. Furman, and Jamshid Mostoufi. UltraSPARC-IIi: Expanding the boundaries of a system on a chip. IEEE Micro, 18(2):14–24, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL hf=

0?target=if(eq(query(%27%
3CFNO%3E+cont+m2014%27),0)
,1,ancestor(ARTICLE,query(
%27%3CFNO%3E+cont+m2014%27)
));   http://dlib.computer.
org/dynaweb/mi/mi1998/@ebt-
link;   http://dlib.computer.
org/mi/books/mi1998/pdf/m2014.
pdf;   http://www.computer.
org/micro/mi1998/m2014abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Nickolls:2010:GCE

[ND10] John Nickolls and William J.

Dally. The GPU computing era. IEEE Micro, 30(2):56– 69, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Nemirovsky:1995:ANI

[Nem95] M. D. Nemirovsky. The architecture of the NS486 integrated processor. In IEEE [IEE95], pages 11–18. ISBN ???? LCCN ????

Ngai:1995:VAF

[Nga95] A. Ngai. VLSI architecture of the I-frame encoder for the MPEG-2 video compression. In IEEE [IEE95], pages 103–110. ISBN ???? LCCN ????

Narad:2000:TNP

[NH00] Chuck Narad and Larry Huston. Tutorial: Network processors. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Nickolls:2003:CLP

[NIJ^{+}03] John Nickolls, L. J. Madar III, Scott Johnson, Viresh Rustagi, Ken Unger, and Mustafiz Choudhury. Calisto: A low-power singlechip multiprocessor communications platform. IEEE Micro, 23(2):29–43, March/ April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2003/pdf/m2029.
pdf;   http://www.computer.
org/micro/mi2003/m2029abs.
htm.

Nguyen:1996:EMS

[NMP^{+}96] L. T. Nguyen, M. Mohamed, H. Park, Y. Pai, R. Wong, A. Qureshi, P. Song, H. D. Truong, and C. Reader. Establish MSP as the standard for media processing. In IEEE [IEE96], pages 217–226. ISBN ???? LCCN ????

Naffziger:2014:HC

[NN14] Samuel Naffziger and Donald Newell. Hot Chips 25. IEEE Micro, 34(2):4–5, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Norrie:2021:DPG

[NPY^{+}21] T. Norrie, N. Patil, D. H. Yoon, G. Kurian, S. Li, J. Laudon, C. Young, N. Jouppi, and D. Patterson. The design process for Google’s training chips: TPUv2 and TPUv3. IEEE Micro, 41(2):56–63, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Naffziger:2015:HC

[NS15] Samuel Naffziger and Guri Sohi. Hot Chips 26. IEEE Micro, 35(2):4–5, March/April 2015.

CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020004-abs.
html.

Normoyle:1997:UI

[Nt97] Kevin Normoyle and the Sabre cats. UltraSparc IIi — A highly integrated 300 MHz 64bit SPARC V9 CPU. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Nanya:1997:TBS

[NTK^{+}97] Takashi Nanya, Akihiro Takamura, Masashi Kuwako, Masashi Imai, Taro Fujii, Motokazu Ozawa, Izumi Fukasaku, Yoichiro Ueno, Fuyuki Okamoto, Hiroki Fujimoto, Osamu Fujita, Masakazu Yamashina, and Masao Fukuma. TITAC-2: A 32-bit scalable-delay-insensitive microprocessor. In IEEE [IEE97], pages 19–32. ISBN ???? LCCN ????

Oehler:1991:IRS

[OB91] Richard R. Oehler and Michael W.

Blasgen. IBM RISC System/ 6000: architecture and performance. IEEE Micro, 11(3):14– 17, 56–62, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

OConnor:2000:NAP

[O’C00a] Mike O’Connor. he network address processor from silicon access: Using wide, high-speed embedded DRAM for fast forwarding table lookup. In IEEE [IEE00], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/index12.html.

OConnor:2000:IAP

[O’C00b] Mike O’Connor. The iFlow address processor: Forwarding table lookups using fast, wide embedded DRAM. In IEEE [IEE00], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/index12.html.

ODonnell:1999:MVM

[O’D99] John Setel O’Donnell. MAP1000A: a 5W, 230MHz VLIW mediaprocessor. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

OConnor:2001:IAP

[OG01] Mike O’Connor and Christopher A. Gomez. The iFlow address processor. IEEE Micro, 21(2):16–23, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

                                        http://dlib.computer.org/
                                        mi/books/mi2001/pdf/m2016.
                                        pdf;   http://www.computer.
                                        org/micro/mi2001/m2016abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Opris:2000:BAF

[OKN^{+}00] Ion Opris, Jonathan Kleks, Yasunori Noguchi, James Castillo, Shiyin Siou, Murty Bhavana, Youichi Nakasone, Shingo Kokudo, and Seiichiro Watanabe. A 12-bit 50Mpixel/s analog front end processor for digital imaging systems. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Orup:1994:SCM

[Oru94] Holger Orup. A 100Kbit/s single chip modular exponentiation processor. In IEEE [IEE94], pages 53–59. ISBN ???? LCCN ????

Owen:2008:NAA

[OS08] Jonathan Owen and Maurice Steinman. Northbridge architecture of AMD’s Griffin microprocessor family. IEEE Micro, 28(2):10–18, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Opris:2001:FAF

[OW01] Ion E. Opris and Seiichiro Watanabe. A fast analog front-end processor for digital imaging systems. IEEE Micro, 21(2):48–54, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2001/pdf/m2048.
pdf;   http://www.computer.
org/micro/mi2001/m2048abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Oberman:1998:ATK

[OWJF98] Stuart Oberman, Fred Weber, Norbert Juffa, and Greg Favor. AMD 3DNow! technology and the K6-2 microprocessor. In IEEE [IEE98], pages 245–254. ISBN ???? LCCN ????

Patterson:2006:RRA

[PAA^{+}06] David Patterson, Arvind, Krste Asanovic, Derek Chiou, James C. Hoe, Christoforos Kozyrakis, Shih-Lien Lu, Mark Oskin, Jan Rabaey, and John Wawrzynek. RAMP: Research Accelerator for Multiple Processors. Web document., 2006. URL http://

                                        www.hotchips.org/archives/
                                        hc18/2_Mon/HC18.S4/HC18.S4T2.
                                        pdf.

Panicacci:1997:ICA

[PAGC^{+}97] R. Panicacci, J. Alphonso-Gibbs, A. Cho, A. Krymski, N. Doudoumopoulos, R. Nixon, S. E. Kemeny, and E. R. Fossum. 1/4-inch CMOS active pixel sensor with smart on-chip functions and full digital interface. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Papworth:1995:OPP

[Pap95] D. Papworth. Optimizing the P6 pipeline. In IEEE [IEE95], pages 31–40. ISBN ???? LCCN ????

Papworth:1996:TPP

[Pap96] David B. Papworth. Tuning the Pentium Pro microarchitecture: Refining a design from the initial goals, performance simulations, trade-offs, and dies to the final product. IEEE Micro, 16(2):8–15, March/April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Papadopoulos:1998:KSX

[Pap98] Greg Papadopoulos. Keynote speaker: XXX. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Park:1998:TCD

[Par98] Hee-Bok Park. Two chipsets for DTV: Compliant with ATSC standard. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Patterson:1996:CID

[PAY96] David Patterson, Tom Anderson, and Kathy Yelick. The case for Intelligent DRAM: IRAM. In IEEE [IEE96], pages 75–94. ISBN ???? LCCN ????

Pennello:1990:CCR

[Pen90] Thomas J. Pennello. Compiler challenges with RISCs. IEEE Micro, 10(1):37–43, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Petrovick:2000:PCI

[Pet00] John Petrovick. POWER4 chip integration. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Pfister:1999:VRT

[Pfi99] Hanspeter Pfister. The vg500 real-time ray-casting ASIC. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Phillip:1998:ATS

[Phi98] Mike Phillip. AltiVec technology: A second generation SIMD microprocessor architecture. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Piazza:1998:IIG

[Pia98] Tom Piazza. Intel i740 graphics accelerator. In IEEE [IEE98], page ?? ISBN ???? LCCN ???? The presentation arrived too late to be included on the Hot Chips 10 CD-ROM.

Papazian:2015:IBS

[PKB^{+}15] Irma Esmer Papazian, Sailesh Kottapalli, Jeff Baxter, Jeff Chamberlain, Geetha Vedaraman, and Brian Morris. Ivy Bridge server: A converged design. IEEE Micro, 35(2):16– 25, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2015/02/mmi2015020016-abs.
html.

Plummer:2000:AIE

[Plu00a] Suzanne Plummer. The Au1000 Internet Edge Processor: A high performance, low power SOC: The first chip in a family of parts from Alchemy Semiconductor, Inc. In IEEE [IEE00], page ?? ISBN ???? LCCN

????  URL http://www.
hotchips.org/index12.html.

Plummer:2000:WIH

[Plu00b] Suzanne Plummer. Wireless Internet: High performance/low power optimization. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Runner-up for best presentation award.

Potter:1997:RHP

[Pot97] Jeff Potter. R3D/100 — 3D high performance chip set. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Pratt:1996:ATT

[Pra96] Yale Pratt. Afternoon tutorial: Toward 10 instructions/ cycle uniprocessors. In IEEE [IEE96], page ?? ISBN ???? LCCN ????

Priem:1990:DGG

[Pri90] Curtis R. Priem. Developing the GX graphics accelerator architecture. IEEE Micro, 10(1): 44–54, January/February 1990. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Pronk:2006:HIN

[Pro06] Ben Pronk. Highly integrated Nexperia PNX8535 hybrid television processor. Web document., 2006. URL http://

www.hotchips.org/archives/

hc18/2_Mon/HC17.S1/HC17.S1T1.
pdf.

Przybylski:1997:MTS

[Prz97] Steven Przybylski. Morning tutorial: Sorting out the new DRAMs. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Pellegrini:2020:ANN

[PSB^{+}20] A. Pellegrini, N. Stephens, M. Bruce, Y. Ishii, J. Pusdesris, A. Raja, C. Abernathy, J. Koppanalil, T. Ringe, A. Tummala, J. Jalal, M. Werkheiser, and A. Kona. The Arm Neoverse N1 platform: Building blocks for the next-gen cloud-to-edge infrastructure SoC. IEEE Micro, 40(2):53–62, March/April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Popescu:1991:MA

[PSS^{+}91] Val Popescu, Merle Schultz, John Spracklen, Gary Gibson, Bruce Lightner, and David Isaman. The Metaflow architecture. IEEE Micro, 11(3):10– 13, 63–73, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Peterson:1991:IML

[PSW91] Craig Peterson, James Sutton, and Paul Wiley. Iwarp — A 100-Mops, LIW microprocessor for multicomputers. IEEE Micro, 11(3):26–29, 81–87, May/ June 1991. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Quach:2000:IPF

[Qua00] Nhon Quach. The Itanium processor features for high availability and reliability. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Rabaey:2006:WHC

[Rab06]  Jan M. Rabaey.  Wireless

in the home: Challenges and opportunities. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/1_Sun/HC18.T2.pdf.

Rathnam:1998:SCD

[Rat98] Selliah Rathnam. A single chip DTV media processor programmable architecture. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Rattner:2006:CCH

[Rat06] Justin Rattner. Cool codes for hot chips. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.Keynote%20One/
HC18.Keynote1.pdf; http://
www.hotchips.org/hc18/docs/
keynote1_hc18.pdf.

Rupley:2019:SMP

[RBGZ19] J. Rupley, B. Burgess, B. Grayson, and G. D. Zuraski. Samsung M3 processor. IEEE Micro, 39(2): 37–44, March/April 2019. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rogenmoser:2013:RTV

[RC13] Robert Rogenmoser and Lawrence T. Clark. Reducing transistor variability for high performance low power chips. IEEE Micro, 33 (2):18–26, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Ruhl:2013:IPW

[RDJ^{+}13] Gregory Ruhl, Saurabh Dighe, Shailendra Jain, Surhud Khare, and Sriram R. Vangal. IA-32 processor with a widevoltage-operating range in 32nm CMOS. IEEE Micro, 33(2): 28–36, March/April 2013. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Renau:2011:HC

[RE11] Jose Renau and Will Eatherton. Hot Chips 22. IEEE Micro, 31 (2):4–5, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Reinhardt:1996:PNH

[Rei96] Dennis Reinhardt. Proceedings notebook for Hot Chips VIII, Stanford University, August 19-20, 1996. In IEEE [IEE96], page ?? ISBN ???? LCCN ????

Rusu:2004:IPH

[RMC04] Stefan Rusu, Harry Muljono, and Brian Cherkauer. Itanium 2 processor 6M: Higher frequency and larger L3 cache. IEEE Micro, 24(2):10–18, March/ April 2004. CODEN IEMIDZ.

ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2004/02/m2010abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2010.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2010.
pdf.

Rotem:2012:PMA

[RNA^{+}12] Efraim Rotem, Alon Naveh, Avinash Ananthakrishnan, Eliezer Weissmann, and Doron Rajwan. Power-management architecture of the Intel microarchitecture code-named Sandy Bridge. IEEE Micro, 32(2): 20–27, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rosenblum:1999:VVP

[Ros99] Mendel Rosenblum. VMware’s Virtual Platform: A virtual machine monitor for commodity PCs. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Radin:1998:IPL

[RSS98] Margaret Jane Radin, William Benjamin Scott, and Luna M. Scott. Intellectual property law as applied to the computer and electronics industries. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Reick:2008:FTD

[RSS^{+}08] Kevin Reick, Pia N. Sanda, Scott Swaney, Jeffrey W.

Kellington, Michael Mack, Michael Floyd, and Daniel Henderson. Fault-tolerant design of the IBM Power6 microprocessor. IEEE Micro, 28(2):30– 38, March/April 2008. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Rubin:1997:DFU

[Rub97] Norm Rubin. Digital FX!32: A utility for fast transparent execution of Win32 x86 applications on Alpha NT. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Raina:2021:BPH

[RY21] P. Raina and C. Young. Best papers from Hot Chips 32. IEEE Micro, 41(2):6, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slegel:1998:IGM

[SAC^{+}98] Timothy Slegel, Robert M. Averill III, Mark A. Check, Bruce C. Giamei, Barry W. Krumm, Christopher A. Krygowski, Wen H. Li, John S. Liptay, John D. MacDougall, Thomas A. McPherson, Jennifer A. Navarro, Eric M. Schwarz, Kevin Shum, and Charles F. Webb. IBM S/390 G5 microprocessor. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Samueli:1999:KBC

[Sam99]  Henry Samueli.  Keynote:

Broadband communications IC’s: Enabling the connected world of the 21st century. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL ftp://www.hotchips.
org//pub/hotc7to11cd/hc99/
hc11_pdf/hc99.keynote.samueli.
txt; http://www.hotchips.
org/HotChips_Talk.pdf.

Samaras:2000:IPC

[Sam00a] Bill Samaras. The Itanium processor cartridge. In IEEE [IEE00], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/index12.html.

Samueli:2000:BR

[Sam00b] Henry Samueli. The broadband revolution. IEEE Micro, 20(2):16–26, March/April 2000. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2000/pdf/m2016.
pdf;   http://www.computer.
org/micro/mi2000/m2016abs.

htm. Presented at Hot Chips 11 Conference, Stanford University, Stanford, California, August 15–17, 1999.

Santhanam:1996:SCA

[San96] Sribalan Santhanam. StrongArm 110: A 160MHz 32b 0.5W CMOS ARM processor. In IEEE [IEE96], pages 119–130. ISBN ???? LCCN ????

Suzuki:1998:VAE

[SANK98] Kazumasa Suzuki, Tomohisa Arai, Kouhei Nadehara, and Ichiro Kuroda. V830R/AV: An embedded multimedia superscalar RISC processor. IEEE Micro, 18(2):36–47, March/ April 1998. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL hf=

0?target=if(eq(query(%27%
3CFNO%3E+cont+m2036%27),0)
,1,ancestor(ARTICLE,query(
%27%3CFNO%3E+cont+m2036%27)
));   http://dlib.computer.
org/dynaweb/mi/mi1998/@ebt-
link;   http://dlib.computer.
org/mi/books/mi1998/pdf/m2036.
pdf;   http://www.computer.
org/micro/mi1998/m2036abs.

htm. Presented at Hot Chips IX, Stanford University, Stanford, California, August 24–26, 1997.

Savell:1998:EDA

[Sav98] Tom Savell. The EMU10K1 digital audio processor. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Snelgrove:2023:SPT

[SB23] Martin Snelgrove and Robert Beachler. speedAI240: a 2petaflop, 30-Teraflops/W atmemory inference acceleration device with 1456 RISC-V cores. IEEE Micro, 43(3):58–63, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stephens:2017:ASV

[SBB^{+}17] Nigel Stephens, Stuart Biles, Matthias Boettcher, Jacob Eapen, Mbou Eyole, Giacomo Gabrielli, Matt Horsnell, Grigorios Magklis, Alejandro Martinez, Nathanael Premillieu, Alastair Reid, Alejandro Rico, and Paul Walker. The ARM scalable vector extension. IEEE Micro, 37(2):26– 39, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020026-abs.
html.

Shum:2013:IZT

[SBJ13] C. Kevin Shum, Fadi Busaba, and Christian Jacobi. IBM zEC12: The third-generation high-frequency mainframe microprocessor. IEEE Micro, 33 (2):38–47, March/April 2013. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Storino:1999:MTB

[SBKK99] Salvatore Storino, John M.

Borkenhagen, Ronald N. Kalla, and Steven R. Kunkel. A multithreaded 64-bit PowerPC commercial RISC processor design. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Smith:1997:WOR

[SBS97] S. Diane Smith, Allen J. Baum, and Alan Jay Smith. Welcome and opening remarks. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Stefan:2006:CFP

[SBS^{+}06] Gheorghe Stefan, Lazar Bivolarski, Anand Sheel, Bogdan Mitu, Tom Thomson, and Dan Tomescu. The CA1024: A fully programmable system-on-chip for cost-effective HDTV media processing. Web document., 2006. URL http://

                                        www.hotchips.org/archives/
                                        hc18/2_Mon/HC18.S5/HC18.S5T2.
                                        pdf.

Schmidt:1991:DSC

[SC91] Ulrich Schmidt and Knut Caesar. Datawave — A single-chip multiprocessor for video applications. IEEE Micro, 11(3):22– 25, 88–94, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Segars:1999:EIA

[Seg99] Simon Segars. Enabling innovation: The ARM9E synthesizable processor family. In IEEE [IEE99], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/hotc11_index.
                                        html.

Sell:2018:XOX

[Sel18] John Sell. The Xbox One X Scorpio Engine. IEEE Micro, 38(2):53–60, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL https://www.
                                        computer.org/csdl/mags/mi/
                                        2018/02/mmi2018020053-abs.
                                        html.

Shreejith:2018:SNI

[SF18] Shanker Shreejith and Suhaib A.

Fahmy. Smart network interfaces for advanced automotive applications. IEEE Micro, 38 (2):72–80, March/April 2018. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2018/02/mmi2018020072-abs.
html.

Sodani:2016:KLS

[SGC^{+}16] Avinash Sodani, Roger Gramunt, Jesus Corbal, Ho-Seop Kim, Krishna Vinod, Sundaram Chinthamani, Steven Hutsell, Rajat Agarwal, and Yen-Chen Liu. Knights Landing: Secondgeneration Intel Xeon Phi product. IEEE Micro, 36(2):34– 46, March/April 2016. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL http://www.
computer.org/csdl/mags/mi/
2016/02/mmi2016020034-abs.
html.

Shah:2012:STD

[SGG^{+}12] Manish Shah, Robert Golla, Gregory Grohoski, Paul Jordan, Jama Barreh, Jeff Brooks, Mark Greenberg, Gideon Levinsky, Mark Luttrell, Christopher Olson, Zeid Samoail, Matt Smittle, and Tom Ziaja. Sparc T4: a dynamically threaded serveron-a-chip. IEEE Micro, 32(2): 8–19, March/April 2012. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Shao:1996:MTJ

[Sha96] Sami Shao. Morning tutorial: Java software secrets. In IEEE [IEE96], page ?? ISBN ???? LCCN ????

Sharangpani:2000:IIP

[Sha00a] Harsh Sharangpani. Intel Itanium processor core. In IEEE [IEE00], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/index12.html.

Sharangpani:2000:IPC

[Sha00b] Harsh Sharangpani. The Itanium processor core. In IEEE [IEE00], page ?? ISBN ????

                                        LCCN ???? URL http://www.
                                        hotchips.org/index12.html.

Runner-up for best presentation award.

Shen:1995:SHS

[She95] G. W. Shen. SPARC64+: HAL’s second generation 64-bit SPARC processor. In IEEE [IEE95], pages 75–86. ISBN ???? LCCN ????

Sheng:1999:SPCa

[She99a] Sam Sheng. Signal processing in communications I: xDSL. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Sheng:1999:SPCb

[She99b] Sam Sheng. Signal processing in communications II: CDMA. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

                                        //www.hotchips.org/hotc11_
                                        index.html.

Shimizu:1998:MDS

[Shi98] Toru Shimizu. M32Rx/D — A single chip microcontroller with A high capacity 4MB internal DRAM. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Saito:1995:MSR

[SHMS95] K. Saito, M. Hashimoto, K. Matsubara, and H. Sawamoto. A 150 MHz superscalar RISC processor with pseudo vector processing feature. In IEEE [IEE95], pages 197–206. ISBN ???? LCCN ????

Subramony:2024:ARS

[SKP24] Mahesh Subramony, David Kramer, and Indrani Paul. AMD Ryzen 7040 series. IEEE Micro, 44(3):18–24, May/June 2024. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sriram:2023:HHS

[SKW^{+}23] Karthik Sriram, Ioannis Karageorgos, Xiayuan Wen, Ján Veselý, Nick Lindsay, Michael Wu, Lenny Khazan, Raghavendra Pradyumna Pothukuchi, Rajit Manohar, and Abhishek Bhattacharjee. HALO: a hardware software co-designed processor for brain computer interfaces. IEEE Micro, 43(3): 64–72, May/June 2023. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sheafor:2000:MCV

[SL00] Steve Sheafor and Cindy Lindsay. Making connections:

Vitesse Network processors (Sitera’s PRISM IQ2000 NPU family): Optimizing architecture for bandwidth and flexibility. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http:/

/www.hotchips.org/index12.
html.

Slaton:1997:SSP

[Sla97] Jeff Slaton. The StrongARM SA-1100: A portable communications microprocessor. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Slavenburg:1999:THS

[SLR^{+}99] Gerrit Slavenburg, Luis Lucas, Thorwald Rabeler, Naeem Maan, Chang-Ming Yang Hani Salloum, Farah Jubran, Mohammed I. Yousuf, Babu Kandimalla, and Muhammad Hafeez. TM-1300 high-speed, low-cost, enhanced PCI, VLIW media processor. In IEEE [IEE99], page ?? ISBN ????

LCCN ???? URL http://www.
hotchips.org/hotc11_index.
html.

Sachs:1991:DIT

[SMHB91] Howard G. Sachs, Harlan McGhan, Lee F. Hanson, and Nathan A. Brookwood. Design and implementation trade-offs in the Clipper C400 architecture. IEEE Micro, 11(3):18–21, 74–80, May/June 1991. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sell:2014:XOS

[SO14] John Sell and Patrick O’Connor.

The Xbox One system on a chip and Kinect sensor. IEEE Micro, 34(2):44–53, March/April 2014. CODEN IEMIDZ. ISSN 0272-1732.

Selvaggi:2009:BMP

[SP09] Richard Selvaggi and Larry Pearlstein. Broadcom mediaDSP: A platform for building programmable multicore video processors. IEEE Micro, 29(2): 30–45, March/April 2009. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Slavenburg:1996:TTP

[SRD96] Gerrit A. Slavenburg, Selliah Rathnam, and Henk Dijkstra. The Trimedia TM-1 PCI VLIW mediaprocessor. In IEEE [IEE96], pages 171–178. ISBN ???? LCCN ????

Sell:2006:GEI

[SS06] John Sell and Alan Jay Smith.

Guest Editors’ introduction: Hot Chips 17. IEEE Micro, 26 (2):8–9, March/April 2006. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Scherer:2022:SIH

[SS22] Alisa Scherer and Guri Sohi.

Special issue on Hot Chips 33. IEEE Micro, 42(3):6, May/ June 2022. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Suggs:2020:AZP

[SSB20] D. Suggs, M. Subramony, and D. Bouvier. The AMD Zen 2 processor. IEEE Micro, 40(2): 45–52, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sugumar:2021:MTN

[SSR21] R. Sugumar, M. Shah, and R. Ramirez. Marvell ThunderX3: Next-generation Armbased server processor. IEEE Micro, 41(2):15–21, March/ April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1990:FAHa

[Ste90a] Bob Stewart, editor. First Annual Hot Chips Symposium, Part 1, volume 10(1) of IEEE Micro. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stewart:1990:FAHb

[Ste90b] Bob Stewart, editor. First Annual Hot Chips Symposium, Part 2, volume 10(3) of IEEE Micro. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, June 1990. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stearns:1995:SSC

[Ste95] C. Stearns. S3 single chip MPEG-1 audio/video decoder.

In IEEE [IEE95], pages 111– 120. ISBN ???? LCCN ????

Stern:2011:SSR

[Ste11]  Richard H. Stern.  Stan-

dardization skullduggery revisited. IEEE Micro, 31(2):96– 99, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Stiles:2019:HSB

[Sti19] D. Stiles. The hardware security behind Azure Sphere. IEEE Micro, 39(2):20–28, March/April 2019. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Sadasivam:2017:IPP

[STKS17] Satish Kumar Sadasivam, Brian W.

Thompto, Ron Kalla, and William J. Starke. IBM Power9 processor architecture. IEEE Micro, 37(2):40– 51, March/April 2017. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020040-abs.
html.

Starke:2021:IPP

[STSM21] W. J. Starke, B. W. Thompto, J. A. Stuecheli, and J. E. Moreira. IBM’s POWER10 processor. IEEE Micro, 41(2):7– 14, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Steffen:2001:TQC

[SVC01] Matthias Steffen, Lieven M. K. Vandersypen, and Isaac L. Chuang. Toward quantum computation: A five-qubit quantum processor. IEEE Micro, 21(2):24–34, March/April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2001/pdf/m2024.
pdf;   http://www.computer.
org/micro/mi2001/m2024abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Tirumalai:1996:PCU

[TGK^{+}96] Partha Tirumalai, Vinod Grover, Xiangyun Kong, Michael Lai, Jian-Zhong Wang, Kurt Goebel, Chris Aoki, Peter Damron, and Krishna Subramanian. A parallelizing compiler for Ultra-SPARC systems. In IEEE [IEE96], pages 67–74. ISBN ???? LCCN ????

Thakkar:1999:ISS

[Tha99] Shreekant (Ticky) Thakkar. The Internet Streaming SIMD extensions. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Talla:2004:APD

[THT^{+}04] Deepu Talla, Ching-Yu Hung, Raj Talluri, Frank Brill, David Smith, David Brier, Bruce Xiong, and Derek Huynh.

Anatomy of a portable digital mediaprocessor. IEEE Micro, 24(2):32–39, March/ April 2004. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL http:

//csdl.computer.org/comp/
mags/mi/2004/02/m2032abs.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2032.
htm;   http://csdl.computer.
org/dl/mags/mi/2004/02/m2032.
pdf.

Taylor:2002:RMC

[TKM^{+}02] Michael Bedford Taylor, Jason Kim, Jason Miller, David Wentzlaff, Fae Ghodrat, Ben Greenwald, Henry Hoffman, Paul Johnson, Jae-Wook Lee, Walter Lee, Albert Ma, Arvind Saraf, Mark Seneski, Nathan Shnidman, Volker Strumpen, Matt Frank, Saman Amarasinghe, and Anant Agarwal. The Raw microprocessor: A computational fabric for software circuits and generalpurpose programs. IEEE Micro, 22(2):25–35, March/ April 2002. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2002/pdf/m2025.
pdf;   http://www.computer.
org/micro/mi2002/m2025abs.
htm.

Tamaki:1999:RPS

[TKS^{+}99] Y. Tamaki, T. Kurihara, K. Shimada, E. Kamada, and T. Shimizu. A RISC processor for SR8000: Accelerating large scale scientific computing with SMP. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/hotc11_index.html.

Tremblay:1996:UFI

[TO96a] Marc Tremblay and J. Michael O’Connor. UltraSparc I: A four-issue processor supporting multimedia: Combining on-chip multimedia instructions with a high-performance, four-issue architecture. IEEE Micro, 16(2): 42–50, March/April 1996. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Tremblay:1996:PHI

[TO96b] Marc Tremblay and Michael O’Connor. PicoJava: A hardware implementation of the Java Virtual Machine. In IEEE [IEE96], pages 131–144. ISBN ???? LCCN ????

Tabatabaei:2010:SMO

[TP10] Sassan Tabatabaei and Aaron Partridge. Silicon MEMS oscillators for high-speed digital systems. IEEE Micro, 30(2): 80–89, March/April 2010. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tremblay:1995:UBS

[Tre95] U. Tremblay. UltraSPARC-I: A 64-bit superscalar processor with multimedia support. In IEEE [IEE95], pages 207–216. ISBN ???? LCCN ????

Trevett:1996:PGD

[Tre96] Neil Trevett. Permedia and GLINT Delta: New generation silicon for 3D graphics. In IEEE [IEE96], pages 275–288. ISBN ???? LCCN ????

Trevett:1997:GGG

[Tre97] Neil Trevett. Glint Gamma: A 3D geometry and lighting processor for the PC. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Trevett:1998:PTG

[Tre98] Neil Trevett. PERMEDIA 3 — A third generation graphics controller for the PC mainstream and DirectX. In IEEE [IEE98], page ?? ISBN ???? LCCN ???? The presentation arrived too late to be included on the Hot Chips 10 CD-ROM.

Tremblay:1999:MAN

[Tre99] Marc Tremblay. MAJC: An architecture for the new millennium. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Truong:1997:VAT

[Tru97] Loc Truong. The VelociTI architecture of the TMS320C6xxx. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Takeo:2006:USH

[TSI06] Akihiko Takeo, Kazuhito Shimomura, and Jun Itoh. The ultra small HDD for the mobile applications. Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.S3/HC18.S3T2.
pdf.

Talpes:2020:CST

[TSV^{+}20] E. Talpes, D. D. Sarma, G. Venkataramanan, P. Bannon, B. McGee, B. Floering, A. Jalote, C. Hsiong, S. Arora, A. Gorti, and G. S. Sachdev. Compute solution for Tesla’s full self-driving computer. IEEE Micro, 40(2):25–35, March/ April 2020. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Tremaine:2001:PIM

[TSW^{+}01] R. Brett Tremaine, T. Basil Smith, Mike Wazlowski, David Har, Kwok-Ken Mak, and Sujith Arramreddy. Pinnacle: IBM MXT in a memory controller chip. IEEE Micro, 21(2):56–68, March/ April 2001. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2001/pdf/m2056.
pdf;   http://www.computer.
org/micro/mi2001/m2056abs.

htm. Presented at Hot Chips 12 Conference, Stanford University, Stanford, California, August 13–15, 2000.

Talpes:2023:MDT

[TSW^{+}23] Emil Talpes, Debjit Das Sarma, Doug Williams, Sahil Arora, Thomas Kunjan, Benjamin Floering, Ankit Jalote, Christopher Hsiong, Chandrasekhar Poorna, Vaidehi Samant, John Sicilia, Anantha Kumar Nivarti, Raghuvir Ramachandran, Tim Fischer, Ben Herzberg, Bill McGee, Ganesh Venkataramanan, and Pete Banon. The microarchitecture of DOJO, Tesla’s exa-scale computer. IEEE Micro, 43(3):31–39, May/ June 2023. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Thekkath:1999:AEE

[TUHwH99] Radhika Thekkath, Mike Uhler, Chandlee Harrell, and Ying wai Ho. An architecture extension for efficient geometry processing. In IEEE [IEE99], page ?? ISBN ???? LCCN ????

URL http://www.hotchips.
org/hotc11_index.html.

Varhol:1994:WAH

[Var94] Peter Varhol. What’s Ahead In Hot Chips. Datamation, 40(22): 73–??, November 15, 1994. CO-DEN DTMNAT. ISSN 0011-6963.

Vasiljevic:2021:CSS

[VBC^{+}21] J. Vasiljevic, L. Bajic, D. Capalija, S. Sokorac, D. Ignjatovic, L. Bajic, M. Trajkovic, I. Hamer, I. Matosevic, A. Cejkov, U. Aydonat, T. Zhou, S. Z. Gilani, A. Paiva, J. Chu, D. Maksimovic, S. A. Chin, Z. Moudallal, A. Rakhmati, S. Nijjar, A. Bhullar, B. Drazic, C. Lee, J. Sun, K. M. Kwong, J. Connolly, M. Dooley, H. Farooq, J. Y. T. Chen, M. Walker, K. Dabiri, K. Mabee, R. S. Lal, N. Rajatheva, R. Retnamma, S. Karodi, D. Rosen, E. Munoz, A. Lewycky, A. Knezevic, R. Kim, A. Rui, A. Drouillard, and D. Thompson. Compute substrate for Software 2.0. IEEE Micro, 41(2):50– 55, March/April 2021. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

vandeWaerdt:2006:PDW

[vdWAB^{+}06] Jan-Willem van de Waerdt, James Akiyama, Bob Brummer, Bill Curtis, Eugene Shteyn, Glen Stone, and Q. Yamada. Panel discussion: Who owns the living room? Web document., 2006. URL http://

www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.dell.pdf;   http://
www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.intel.pdf;   http://
www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.microsoft.pdf;   http:
//www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.Q@KIT.pdf;   http://
www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.samsung.pdf; http://
www.hotchips.org/archives/
hc18/2_Mon/HC18.Panel/HC18.
Panel.sony.pdf.

vanEijndhoven:1998:NMI

[vES98] J. T. J. van Eijndhoven and F. W. Sijstermans. Novel multimedia instruction capabilities in VLIW media processors. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Vinge:2007:KTN

[Vin07]  Vernor Vinge.  Keynote

I: True names and rainbows end. Web document., July/August 2007. URL

http://www-rohan.sdsu.edu/
faculty/vinge/hotchips/index.
htm.

Vissers:1999:VAA

[Vis99] Kees A. Vissers. Video algorithms and architectures. In IEEE [IEE99], page ?? ISBN ???? LCCN ???? URL http:

//www.hotchips.org/hotc11_
index.html.

Viterbi:2000:KAE

[Vit00] Andrew J. Viterbi. Keynote address: Enabling the next generation of wireless devices. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http:/

/www.hotchips.org/index12.
html.

Volos:2017:FCS

[VJFG17] Stavros Volos, Djordje Jevdjic, Babak Falsafi, and Boris Grot. Fat caches for scaleout servers. IEEE Micro, 37 (2):90–103, March/April 2017. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

tronic). URL https://www.
computer.org/csdl/mags/mi/
2017/02/mmi2017020090-abs.
html.

Wade:2020:TCT

[WAA^{+}20] M. Wade, E. Anderson, S. Ardalan, P. Bhargava, S. Buchbinder, M. L.Davenport, J. Fini, H. Lu, C. Li, R. Meade, C. Ramamurthy, M. Rust, F. Sedgwick, V. Stojanovic, D. Van Orden, C. Zhang, C. Sun, S. Y. Shumarayev, C. O’Keeffe, T. T. Hoang, D. Kehlet, R. V. Mahajan, M. T. Guzy, A. Chan, and T. Tran. TeraPHY: A chiplet technology for low-power, highbandwidth in-package optical I/O. IEEE Micro, 40(2):63– 71, March/April 2020. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Warneke:2000:PSD

[WAP00] Brett Warneke, Bryan Atwood, and Kristofer S. J. Pister. Preliminary smart dust mote. In IEEE [IEE00], page ?? ISBN ???? LCCN ???? URL http:/

/www.hotchips.org/index12.
html.

Warton:1997:PSI

[War97] John Warton. Panel session: If I were defining ‘Merced’. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Wang:1995:DMP

[WBC^{+}95] Karl Wang, Chris Bryant, Mike Carlson, Tom Elmer, Adrian Harris, Michael Garcia, C. S. Hui, C. K. Leung, Brian Reynolds, Raymond Tang, Laura Weber, Jim Wenzel, Glen Wilson, and Mike Becker. Designing the MPC105 PCI bridge/memory controller. IEEE Micro, 15(2): 44–49, March/April 1995. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VI, Stanford University, CA, August 14–16, 1994.

Wharton:1996:EPS

[WBC^{+}96] John Wharton, John Banning, Brian Case, David S. Hardin, Martin Hopkins, John Novitsky, and Marc Tremblay. Evening panel session, Lagunita Court: Software or silicon: What’s the best route to Java? In IEEE [IEE96], pages 145–146. ISBN ???? LCCN ????

Wawrzynek:2003:GEI

[WD03] John Wawrzynek and Keith Diefendorff. Guest Editors’ introduction: Hot Chips 14— innovation in the face of uncertain economics. IEEE Micro, 23(2):8–11, March/April 2003. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). URL

http://dlib.computer.org/
mi/books/mi2003/pdf/m2008.
pdf.

Webb:2008:IZN

[Web08] Charles F. Webb. IBM z10:

The next-generation mainframe microprocessor. IEEE Micro, 28(2):19–29, March/April 2008. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Weiser:1996:TCP

[Wei96] Uri Weiser. Trade-off considerations and performance of Intel’s MMX technology. In IEEE [IEE96], pages 147–156. ISBN ???? LCCN ????

Wei:2000:SSS

[Wei00] James Y. Wei. A synchronous serial 64 × 64 self routing crossbar chip for multi-terabit switch fabrics. In IEEE [IEE00], page ?? ISBN ???? LCCN ????

                                        URL http://www.hotchips.
                                        org/index12.html.  Check:

was this Hot Chips 11 (1999) or 12 (2000): Web site is confused??

Wharton:1998:SPC

[Wha98] John Wharton. Session 6 panel: Confronting the Microsoft challenge. In IEEE [IEE98], page ?? ISBN ???? LCCN ????

Wittenbrink:2011:FGG

[WKP11] Craig M. Wittenbrink, Emmett Kilgariff, and Arjun Prabhu. Fermi GF100 GPU architecture. IEEE Micro, 31(2):50– 59, March/April 2011. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Watts:2009:VPB

[WMSH09] Lloyd Watts, Dana Massie, Allen Sansano, and Jim Huey. Voice processors based on the human hearing system. IEEE Micro, 29(2):54–63, March/ April 2009. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

Wong:2003:MTC

[Won03] Alfred K. Wong. Microlithography: Trends, challenges, solutions, and their impact on design. IEEE Micro, 23(2):12– 21, March/April 2003. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).

URL http://dlib.computer.
org/mi/books/mi2003/pdf/m2012.
pdf;   http://www.computer.
org/micro/mi2003/m2012abs.
htm.

Winterbottom:1997:DIV

[WP97] Phil Winterbottom and Rob Pike. The design of the Inferno virtual machine. In IEEE [IEE97], page ?? ISBN ???? LCCN ????

Yeager:1996:MRS

[Yea96] Kenneth C. Yeager. The MIPS R10000 superscalar microprocessor: Emphasizing concurrency and latency-hiding techniques to efficiently run large, real-world applications. IEEE Micro, 16(2):28–40, March/ April 1996. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic). Presented at Hot Chips VII, Stanford University, Stanford, California, August 1995.

Yeh:2006:LPH

[Yeh06]  Tse-Yu Yeh.  The low-

power high-performance architecture of the PWRficient processor family. Web document., 2006. URL http://

www.hotchips.org/archives/

                                        hc18/2_Mon/HC18.S2/HC18.S2T1.
                                        pdf.

Yoshida:2015:SXF

[YHT^{+}15] Toshio Yoshida, Mikio Hondou, Takekazu Tabata, Ryuji Kan, Naohiro Kiyota, Hiroyuki Kojima, Koji Hosoe, and Hiroshi Okano. Sparc64 XIfx: Fujitsu’s next-generation processor for high-performance computing. IEEE Micro, 35(2): 6–14, March/April 2015. CO-DEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (elec-

                                        tronic). URL http://www.
                                        computer.org/csdl/mags/mi/
                                        2015/02/mmi2015020006-abs.
                                        html.

Yoshikawa:2006:IHA

[YYA06] Takashi Yoshikawa, Yutaka Yamada, and Shigehiro Asano. An implementation of hardware accelerator using dynamically reconfigurable architecture. Web document., 2006. URL http://

                                        www.hotchips.org/archives/
                                        hc18/2_Mon/HC18.S4/HC18.S4T3.
                                        pdf.

Zaruba:2021:MCR

[ZSB21] F. Zaruba, F. Schuiki, and L. Benini. Manticore: A 4096core RISC-V chiplet architecture for ultraefficient floatingpoint computing. IEEE Micro, 41(2):36–42, March/April 2021. CODEN IEMIDZ. ISSN 0272-1732 (print), 1937-4143 (electronic).